### v1

# T-FIX: Text-Based Explanations with Features Interpretable to eXperts

Abstract
As LLMs are deployed in knowledge‑intensive settings (e.g., surgery, astronomy, therapy), users expect not just answers, but also meaningful explanations for those answers. In these settings, users are often domain experts (e.g., doctors, astrophysicists, psychologists) who require explanations that reflect expert-level reasoning. However, current evaluation schemes primarily emphasize plausibility or internal faithfulness of the explanation, which fail to capture whether the content of the explanation truly aligns with expert intuition. We formalize expert alignment as a criterion for evaluating explanations with T‑FIX, a benchmark spanning seven knowledge-intensive domains. In collaboration with domain experts, we develop novel metrics to measure the alignment of LLM explanations with expert judgment. [^1] https://github.com/BrachioLab/FIX-2
T-FIX: Text-Based Explanations with Features Interpretable to eXperts
[…]
1 Introduction
Figure 1: Current evaluations of LLM explanations typically consider two dimensions: plausibility — whether the reasoning is logically coherent and faithfulness — whether it reflects the model’s true decision process. We introduce a third, orthogonal dimension: expert alignment — whether the LLM reasons as a domain expert would. For instance, an LLM may correctly predict sepsis risk with a plausible and faithful explanation, yet because it relies on features clinicians rarely use, its expert alignment is low.
Figure 2: An overview of the T-FIX construction process. For each dataset, we first establish expert alignment criteria – features deemed important by domain experts for a specific task – through collaboration with these experts and LLM-based deep research tools. These criteria form the basis of the T-FIX evaluation pipeline, which processes an LLM-generated explanation to output an expert alignment score. A high score suggests the explanation reflects reasoning aligned with domain experts (i.e., the LLM “thinks like an expert”), while a low score indicates the explanation may rely on aspects that experts would deem irrelevant.
LLMs are increasingly used for domain-specific tasks that require substantial knowledge from specialized fields — the adoption of LLMs is actively being explored in high-stakes environments like operating rooms, astronomical observatories, and therapy clinics pressman2024clinical; fouesneau2024rolelargelanguagemodels; stade2024large. For LLMs to be trustworthy in these critical applications, users require not only correct answers but also good explanations rudin2019stop; pedreschi2019meaningful.
However, what constitutes a “good explanation” depends on the target audience ribera2019can; sokol2020one. As LLMs are used for specialized tasks like surgical assistance or supernova analysis, the primary users are often domain experts, like doctors or astrophysicists. Consequently, a “good explanation” in these specialized contexts must offer insights that are valuable and interpretable to these domain experts.
Existing evaluations of LLM explanations focus primarily on two dimensions: (1) plausibility, which assesses whether an explanation appears intuitively reasonable to an average human observer; and (2) faithfulness, which measures whether the explanation accurately reflects the LLM’s internal reasoning process zhou2021evaluating; agarwal2024faithfulness; parcalabescu2023measuring.
While these dimensions are necessary, they are insufficient for knowledge-intensive domains. Experts do not merely ask whether an explanation seems plausible or mirrors the model’s internal reasoning; they require that the reasoning itself aligns with domain-validated principles wang2021explanations. In such settings, the key question is whether LLMs consider aspects of the input that experts themselves deem scientifically critical.
This limitation is illustrated in Figure 1: an LLM may generate a correct answer with a plausible and faithful explanation, yet still rely on features irrelevant to the expert’s reasoning framework. Such reasoning, though logically coherent, fails to meet expert standards. To address this, we introduce a third evaluation dimension: expert alignment.
Expert alignment assesses whether an explanation emphasizes the criteria a domain expert would use when making the same prediction. It is orthogonal to plausibility and faithfulness: while plausibility evaluates logical coherence and faithfulness evaluates causal accuracy, expert alignment tests for epistemic validity — that the explanation is grounded in established scientific knowledge.
While prior work has examined expert reasoning via structured feature attribution jin2024fix, such methods are largely confined to classification settings. In contrast, modern LLMs produce free-form textual explanations not directly linked to explicit feature groups. To our knowledge, no existing benchmark evaluates the alignment of such free-form explanations.
To fill this gap, we introduce the T-FIX benchmark: a suite of datasets spanning seven domains with an accompanying evaluation framework. Developed with domain experts, T-FIX quantifies the extent to which LLM explanations align with scientifically-grounded expert reasoning.
Our contributions are as follows:
  • We introduce expert alignment as a desired attribute of LLM-generated explanations and create T-FIX, the first benchmark designed to evaluate this.
  • We release a pipeline to evaluate how well any LLM “thinks like an expert,” designed to be easily extendable to new domains.
  • We show that while LLMs perform better when reasoning over multiple expert criteria, they still struggle to align their explanations with expert intuition, revealing a key direction for future improvement.
2 Expert Alignment Criteria
Figure 3: Our T-FIX pipeline. To evaluate an LLM-generated explanation, we first decompose it into atomic claims. Next, we filter out irrelevant claims, such as unsupported or speculative statements. Each remaining claim is then scored against the domain-specific expert alignment criteria: a score of “complete” indicates perfect overlap with at least one criterion, while “none” indicates no overlap. Filtered-out claims are automatically assigned a score of “none”. We compute the final expert-alignment score for the explanation by averaging across all claim scores.
The development of the T-FIX benchmark was a highly collaborative and interdisciplinary process. For each of our seven domains (see Figure 4), our first step was to identify the expert criteria most relevant to making a prediction, detailed in the left of Figure 2.
When answering knowledge-intensive questions like “Will this patient develop sepsis in the next 12 hours?” or “What kind of supernova produced these wavelengths?”, doctors and astrophysicists rely on domain-specific heuristics, prioritizing certain features over others based on training and experience. For instance, in sepsis classification, an experienced clinician would typically emphasize features like advanced age and hypotension, while assigning lower importance to signals like glucose levels or patient demeanor, which are less directly indicative of sepsis risk.
Thus, an LLM that makes the correct prediction by attending to age and hypotension is more expert-aligned than one that arrives at the same answer by focusing on glucose and demeanor. We define the subset of features that experts prioritize when performing a task as expert alignment criteria.
Step 1: Surveying the field.
To seed our initial list of expert criteria, we prompt OpenAI’s o3 model to perform a comprehensive literature review of the relevant field. Each prompt includes a task description, example input-output pairs from the dataset, and instructions to generate a list of criteria considered important for performing the task – accompanied by reputable citations.
We begin with this deep research approach to avoid over-reliance on any single expert’s perspective. Our goal is to synthesize insights from a broad array of books, journals, and academic publications to produce as comprehensive a list as possible.
Step 2: Iteration with domain experts.
To validate and improve the output from Step 1, we present the preliminary criteria list to a domain expert (see Figure 4 for details on each expert per domain). We ask the expert to (1) remove any incorrect or irrelevant criteria, (2) add any important ones that were missed, and (3) ensure that the list reflects a consensus that their peers would accept. The expert then refines the list until it accurately captures the field’s knowledge.
An example criterion for sepsis classification is as follows: Advanced age (over 65 years) markedly increases susceptibility to rapid sepsis progression and higher mortality after infection.
All Deep Research prompt templates and final expert alignment criteria lists for all domains are available in our GitHub repository.
3 T-FIX Pipeline
Figure 4: Overview of datasets and domains in T-FIX. We evaluate LLM expert alignment across seven diverse domains, spanning cosmology, psychology, and medicine. For each dataset, we highlight the motivating task, input–output format, representative example, and the expert responsible for validating alignment criteria. The final row summarizes the expert alignment criteria used for scoring explanations in each domain. The column colors reflect dataset modality: blue indicates vision, yellow indicates language, and pink indicates time-series.
LLM-generated explanations contain a mix of reasoning steps – some aligned with expert judgment, and others based on irrelevant information.
To systematically evaluate such complex explanations, we first break them down into atomic claims, or standalone “features” that can be individually assessed for expert alignment. By scoring each feature separately and then aggregating these scores, we can compute an overall expert alignment score for the full explanation. See Figure 3 for an example of this multi-step process.
Our T-FIX pipeline, built using GPT-4o, for evaluating expert alignment consists of three components:
  1. 1.
    Claim Extraction: Decomposing a free-form explanation into standalone, atomic claims.
  2. 2.
    Relevancy Filtering: Removing claims that are unsupported, speculative, or otherwise irrelevant to the model’s prediction.
  3. 3.
    Alignment Scoring: Measuring the degree of overlap between each remaining claim and domain expert criteria.
3.1 Stage 1: Atomic Claim Extraction
Given a free-form text explanation accompanying an LLM’s prediction, our first goal is to identify and extract the distinct reasoning steps, i.e. “features”, used by the LLM. We achieve this by decomposing the explanation into atomic claims.
An atomic claim is defined as a self-contained, indivisible statement that conveys a single verifiable fact, and can be fully understood without reference to the surrounding context.
To extract atomic claims, we adapt prompting techniques from the claim decomposition literature wanner2024closer; gunjal2024molecularfactsdesideratadecontextualization and prompt GPT-4o to transform a free-form explanation into a list of fully decontextualized atomic claims. We treat each claim as representing a single “feature” in the LLM’s explanation.
3.2 Stage 2: Relevancy Filtering
Not all extracted claims contribute meaningfully to expert reasoning. Some may be unsupported (i.e., references to content not present in the input), speculative (i.e., unfounded hypotheses), or otherwise irrelevant (e.g., repeating the model’s final prediction or citing unrelated information).
Given domain experts’ preference for succinct, informative explanations, we prompt GPT-4o to remove such noisy claims by evaluating each atomic claim based on the original input. A claim is retained if it satisfies the following two criteria:
(1) Clearly grounded in and supported by the input (i.e., not unfounded or speculative)
(2) Directly contributes to explaining why the model made its prediction.
On average, 72% of the claims generated in Stage 1 pass this relevancy filter and are carried forward for alignment scoring.
3.3 Stage 3: Alignment Scoring
In the final stage of our pipeline, we evaluate each retained atomic claim by comparing it to the domain-specific expert alignment criteria (see Section 2). This step quantifies how the reasoning in the LLM’s explanation reflects expert judgment.
Given an atomic claim and a list of expert criteria, we prompt GPT-4o to measure the claim’s expert alignment in two steps:
  1. 1.
    Identify the most aligned expert criterion. The model selects the criterion whose focus and intent best match the core idea of the atomic claim. The model may also indicate that no criteria align with the claim.
  2. 2.
    Label claim alignment. The model scores how well the claim aligns with the chosen criterion: complete, partial, or none. See Table˜1 for details.
Label Meaning
none The claim references something unrelated to the expert category, or misinterprets the category’s meaning.
partial The claim partially refers to the expert category but lacks key details, uses vague language, is overly general, or contains noise.
complete The claim is specific, directly relevant, and fully captures the meaning and intent of the expert category.
Table 1: Interpretation of alignment labels in scoring atomic claims against expert criteria.
For example, consider the expert criterion for sepsis classification: Advanced age (over 65 years). The claim ‘‘The patient is at risk as they are 72 years old’’ would be labeled complete, as it directly and fully supports the criterion. In contrast, the claim ‘‘The patient is at risk as they are 37’’ may be labeled partial: while it discusses patient age, the specific value does not align with the expert threshold for elevated risk. In contrast, the claim ‘‘The patient is NOT at risk as they are 37’’ would be labeled complete.
Examples of claims with high and low alignment for each domain, along with rationale for why those scores were assigned, are provided in Table˜A4.
3.4 Final Aggregation
We assign an alignment label of none to the claims that were filtered out or did not align with any criteria. This ensures LLM-generated explanations are penalized for unsupported or speculative statements, irrelevant information, and misaligned reasoning. Then, we map alignment labels to numerical scores as follows: complete $\rightarrow 1$, partial $\rightarrow 0.5$, and none $\rightarrow 0$. We average these scores across all claims to produce a final expert alignment score for the explanation.
The prompts for all three stages can be found in Appendix˜B and in our Github repository.
4 Pipeline Validation
Given our pipeline relies on multiple curated GPT-4o prompts, we want to ensure that the extracted and filtered claims are accurate, and that the final alignment scores match domain expert intuition. To do this, we conduct a quantitative evaluation via annotation, and a qualitative evaluation via domain expert review.
Annotation study.
To quantitatively validate the outputs at each stage, we manually annotate 35 examples (5 per domain), covering 295 extracted claims and 211 aligned claims. Six annotators participate, with two annotators per example. [^2] Annotators are PhD students in machine learning at an American university with experience evaluating LLM outputs.
Annotators evaluate each stage using a unified labeling scheme: (A) fully accurate, (B) partially accurate, or (C) incorrect. For atomic claim extraction, they review the explanation and its extracted claims, labeling outputs as (A) correct decomposition, (B) 1–2 claims missing/incorrect, or (C) 3+ claims missing/incorrect. For relevancy filtering, they assess whether claims were appropriately kept or removed, labeling results as (A) all correct, (B) 1–2 incorrect, or (C) 3+ incorrect. For expert alignment scoring, annotators review the filtered claims and expert categories, labeling each as (A) matches expert judgment, (B) borderline, or (C) does not match expert judgment.
These labels are then mapped to accuracy scores of 1.0, 0.5, and 0.0, respectively. Table˜2 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX domains, indicating moderate-to-substantial agreement and supporting the reliability of our pipeline. Domain-level metrics appear in Table˜A3.
Domain expert interview.
To qualitatively validate the expert alignment scores in Stage 3 of our pipeline, we interview each domain expert and have them analyze 10-20 claims for their dataset, discussing noticed successes and failures.
Across domains, experts largely agreed with the LLM alignment labels, confirming that the evaluation framework reliably captures expert reasoning. Agreement was highest in domains with well-defined interpretive criteria (e.g., physics and psychology) and somewhat lower in medicine, where decomposing multi-symptom medical statements occasionally caused partial rather than full alignment. Overall, expert validation supports the interpretability and trustworthiness of the T-FIX evaluation pipeline. We provide a summary of each domain expert’s interview in Appendix˜E.
Pipeline Stage $\mathcal{N}$ Accuracy Cohen’s $\kappa$
Claim Extraction 35 0.943 0.717
Relevancy Filtering 295 0.871 0.402
Expert Alignment 211 0.923 0.405
Table 2: Pipeline validation: Average accuracy for $\mathcal{N}$ samples across all T-FIX domains and annotator agreement – Cohen’s $\kappa$ for each stage in our pipeline. Domain-specific statistics are provided in Table˜A3.
5 Included Datasets
Cosmology Psychology Medicine
Baseline Mass Maps Supernova Politeness Emotion Cholecystectomy Cardiac Sepsis
GPT-4o (openai2024gpt4ocard)
Vanilla 0.334 0.726 0.564 0.494 0.236 0.427 0.535
Chain-of-Thought 0.320 0.697 0.534 0.514 0.278 0.472 0.541
Socratic Prompting 0.329 0.708 0.536 0.501 0.299 0.466 0.547
SubQ Decomposition 0.279 0.719 0.523 0.439 0.294 0.435 0.553
o1 (openai2024openaio1card)
Vanilla 0.528 0.570 0.560 0.521 0.370 0.446 0.525
Chain-of-Thought 0.501 0.552 0.562 0.555 0.387 0.428 0.560
Socratic Prompting 0.428 0.575 0.502 0.397 0.371 0.370 0.555
SubQ Decomposition 0.430 0.582 0.482 0.480 0.328 0.398 0.574
Gemini-2.0-Flash (google2024gemini2flash)
Vanilla 0.454 0.650 0.552 0.495 0.335 0.445 0.544
Chain-of-Thought 0.442 0.666 0.505 0.489 0.312 0.465 0.539
Socratic Prompting 0.223 0.666 0.490 0.470 0.323 0.409 0.560
SubQ Decomposition 0.323 0.660 0.509 0.449 0.325 0.408 0.557
Claude-3.5-Sonnet (anthropic2025claude35)
Vanilla 0.622 0.581 0.558 0.533 0.211 0.459 0.564
Chain-of-Thought 0.605 0.604 0.565 0.493 0.228 0.440 0.563
Socratic Prompting 0.616 0.613 0.529 0.474 0.235 0.448 0.557
SubQ Decomposition 0.532 0.594 0.554 0.503 0.221 0.460 0.530
Table 3: Evaluating top LLMs on T-FIX. We report the average expert alignment score across all examples in the dataset. Corresponding accuracies are in Table˜A2 and baseline prompting strategies are described in Section 6. In general, more sophisticated prompting strategies do not necessarily improve expert alignment of explanation.
T-FIX contains seven open-source datasets, spanning the fields of cosmology, psychology, and medicine. To assess LLM explanations across multiple modalities, we include text, vision, and time-series datasets. We select these seven datasets due to the availability of domain experts willing to work with us for these tasks.
As running T-FIX requires querying LLMs, many of which follow a pay-as-you-go API structure, we cap the benchmark at 700 (100 per dataset) in order for T-FIX to be accessible to as many researchers as possible.
We select a subset of 100 examples from the test set of each open-source dataset in T-FIX, and balance this sampling across classes when possible. We provide an overview of the included open-source datasets in Figure 4. See Appendix˜D for additional details about the motivation, task, and prompting procedure for each dataset.
6 Experiments
After building a pipeline to evaluate the expert alignment of an LLM explanation, we evaluate a suite of today’s top LLMs on T-FIX to determine how expert-aligned these models are on domain-specific tasks. We use the following prompting techniques as baselines to generate explanations for each dataset in T-FIX.
  1. 1.
    Vanilla: The LLM is prompted to generate an explanation along with its answer, without any additional guidance or reasoning structure.
  2. 2.
    Chain-of-Thought (CoT): The LLM is prompted to reason step-by-step through intermediate steps before answering, supporting more accurate responses on multi-step tasks.
  3. 3.
    Socratic Prompting: The LLM is instructed to question its own reasoning, encouraging reflection and reconsideration of uncertainties or assumptions.
  4. 4.
    Subquestion Decomposition: The LLM is guided to break down a complex task into simpler subquestions, answer them individually, and then synthesize a final response.
Domain-specific prompts are in Appendix˜D, with templates for the above prompting strategies in Figure˜A5. Results for GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet  [^3] We only select LLMs with vision support and context windows long enough to accommodate our time-series datasets. All models are accessed in September 2025. are in Table˜3.
Role of LLM in pipeline.
Because our prompts are tightly constrained and refined through multiple iterations, the evaluation pipeline is largely robust to the choice of LLM we use. We validate this by comparing results across three evaluator models — GPT-4o, Gemini-2.0-Flash, and Qwen2.5-VL-7B-Instruct — on one dataset per domain. Despite minor variations, rankings and alignment score ranges remain consistent, confirming that the pipeline is robust to evaluator choice. Further details are provided in Appendix˜C.
7 Analysis
In this section, we analyze how LLMs distribute reasoning across expert criteria and whether higher task accuracy indicates better expert alignment.
Figure 5: Shannon Entropy of expert alignment criteria for GPT-4o. For each prompting baseline, we show coverage of each domain’s explanations across all expert criteria – a high value indicates the LLM considers many criteria across examples, while a low value indicates the LLM focuses on the same criteria repeatedly.
Figure 6: Expert Alignment vs. Accuracy Correlation Heatmap, averaged across GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet. Blue indicates positive correlation, red is negative, gray is no correlation.
7.1 Coverage of Expert Alignment Criteria
Section 3 describes our pipeline for measuring the proportion of expert-aligned claims in LLM explanations. We now examine a complementary question: How many expert alignment criteria does an LLM consider across its explanations?
A single gold-standard explanation rarely requires reasoning over all expert criteria; most high-quality explanations reference only 3–5. Thus, assessing coverage at the question level is not meaningful. Instead, we analyze coverage at the dataset level – whether different prompting strategies lead to a broader utilization of expert criteria across all questions within a domain.
Figure 5 presents the Shannon entropy of GPT-4o’s covered expert criteria in each domain. We observe a correlation between entropy and performance: domains where GPT-4o underperforms (e.g., Cholecystectomy, Supernova) show lower entropy, indicating limited criteria coverage. In contrast, well-performing domains (e.g., Politeness, Sepsis) exhibit more uniform coverage, equally taking into account all expert criteria.
This suggests that LLMs that reason uniformly over expert alignment criteria perform better – a promising insight for future work in prompting or training models to incorporate a broader range of expert reasoning.
7.2 Expert Alignment vs. Accuracy
T-FIX focuses on evaluating explanation quality, but we are also interested in understanding the relationship between expert alignment and prediction accuracy. Specifically, we ask: Does higher answer accuracy correspond to stronger expert alignment?
Figure 6 shows the Pearson correlation of expert alignment (see Table˜3) with accuracy (see Table˜A2) for each domain, averaged across models. In some domains with higher performance, like Cholecystectomy and Emotion, we do observe higher expert alignment as well. However, the overall correlation is weak across domains.
The heatmap suggests today’s high-performing LLMs do not appear to rely on expert reasoning. Future research is needed to explore whether aligning model reasoning with expert criteria – via training objectives or prompting – can improve downstream performance.
8 Related Work
Evaluating LLM explanations.
Common explanation methods for LLMs include feature attribution (e.g., LIME, SHAP ribeiro2016lime; lundberg2017shap), counterfactuals, and self-generated explanations im2023evaluatingutilitymodelexplanations; zhao2023explainabilitylargelanguagemodels. Some models are also trained to produce human-readable justifications camburu2018esnli. To assess explanation quality and utility, recent work highlights criteria such as faithfulness (alignment with the model’s reasoning) and plausibility (how convincing it is to humans) jacovi2020faithfulness; zhou2021evaluating; agarwal2024faithfulness. Human studies show mixed outcomes: explanations sometimes aid understanding hase2020evaluating; bansal2021whole, but can also offer little value or cause over-trust wang2023fair. A promising alternative is to use LLMs as automatic judges of explanation quality zheng2023judging; chen2024judge, providing a scalable substitute for expensive human evaluation; we adopt this approach in T-FIX.
Domain & expert alignment.
Concept-based models constrain parts of the network to predict high-level, human-defined concepts, enabling incorporation of domain knowledge into final predictions koh2020concept. Extensions of concept bottlenecks and related methods aim to align latent representations with semantically meaningful features kim2018tcav; chen2020concept; ghorbani2019ace, potentially grouped for expert interpretability jin2024fix. In NLP, integrating human knowledge has included collecting human-written explanation datasets to train models camburu2018esnli and using learned explanations to guide predictions bhatt2020explainable. To our knowledge, no prior work explicitly evaluates text explanations for expert alignment like T-FIX.
9 Conclusion
We introduce T-FIX, the first benchmark designed to evaluate LLM explanations for expert alignment across seven knowledge-intensive domains. Our analysis reveals that today’s models struggle to generate explanations that experts would rely on, highlighting a critical area for improvement.
Future work may include exploring instruction-tuning LLMs to generate explanations with strong expert alignment, extending T-FIX to additional domains, and Human-Computer Interaction studies exploring how expert-aligned explanations affect real-world decision-making by practitioners.
Limitations
As with any LLM-based system, the quality of the outputs is dependent on the input prompt. T-FIX is no exception – though we spend a significant amount of time analyzing outputs and prompt iterating, we do a finite amount of prompt iteration. There is a chance our benchmark could be marginally improved with additional prompt iteration. We hope the issue of prompt dependency diminishes with future models that are more robust and less susceptible to tiny prompt ablations.
While our evaluation pipeline currently uses GPT-4o for scoring, it is model-agnostic by design, and we encourage future work to apply or adapt the pipeline with other LLMs to improve robustness and reduce evaluator-model entanglement.
For pipeline validation, we conduct a user study where we annotate 35 examples. Though the annotation results on this subset suggest our pipeline is accurate, this work could have benefited from a larger and more robust annotation study. Future work should also involve domain experts vetting the pipeline in addition to recruited annotators.
In addition, we only have one expert to validate the expert alignment criteria for each domain. Although our usage of a deep research LLM minimizes over-reliance on a single domain expert, multiple experts would have been better to create the expert criteria. We were constrained by the availability of domain experts.
Our experiments focus on a set of four models and four prompting strategies, and including additional models and strategies could provide a more comprehensive set of baseline results. Though many other high-performing LLMs and prompting techniques exist as of September 2025, we are conscious of budget and the environmental impact of running multiple experiments using T-FIX.
Ethical Considerations
Using LLMs in the domains we describe in T-FIX, especially those relating to medicine, poses a unique set of risks and challenges. We do not advocate that LLMs should replace domain experts in these tasks; rather, T-FIX should serve as a step towards experts being able to use LLMs in a reliable and trustworthy way.
Additionally, LLMs are constantly changing, especially those that are company-owned and not open-source. This poses potential issues relating to the reproducibility of our baseline results as time progresses and advances are made.
Lastly, nearly all LLMs contain biases – some harmful – that may propagate up in a system built off of these models. All users of T-FIX must be conscious of this risk.
Evaluation Model
GPT-4o
Gemini-2.0-Flash
Qwen2.5-VL-7B-Instruct
(Qwen2.5-VL)
Mass Maps
GPT-4o 0.334 0.394 0.507
o1 0.528 0.581 0.596
Gemini-2.0-Flash 0.454 0.483 0.431
Claude-3.5-Sonnet 0.622 0.692 0.651
Emotion
GPT-4o 0.494 0.595 0.579
o1 0.521 0.632 0.567
Gemini-2.0-Flash 0.495 0.621 0.567
Claude-3.5-Sonnet 0.533 0.610 0.565
Cholecystectomy
GPT-4o 0.236 0.000 0.145
o1 0.370 0.000 0.234
Gemini-2.0-Flash 0.335 0.000 0.153
Claude-3.5-Sonnet 0.211 0.000 0.146
Table A1: Model ablation study for evaluation pipeline. Across domains, model rankings are largely consistent across evaluators, indicating the robustness of our evaluation pipeline. Comparison of base models (explanation generator LLMs) across evaluator models (pipeline LLMs) for one dataset per domain. Best per evaluator is in bold; second-best is underlined.
Cosmology Psychology Medicine
Baseline Mass Maps Supernova Politeness Emotion Cholecystectomy Cardiac Sepsis
GPT-4o
Vanilla 0.039 0.103 0.916 0.259 0.075 0.567 0.657
Chain-of-Thought 0.044 0.093 0.824 0.286 0.103 0.460 0.714
Socratic Prompting 0.044 0.127 0.829 0.277 0.115 0.462 0.657
SubQ Decomposition 0.049 0.118 0.837 0.304 0.115 0.485 0.657
o1
Vanilla 0.044 0.170 0.784 0.304 0.194 0.656 0.752
Chain-of-Thought 0.045 0.146 0.818 0.339 0.177 0.685 0.750
Socratic Prompting 0.042 0.155 0.793 0.348 0.155 0.646 0.755
SubQ Decomposition 0.044 0.147 0.818 0.321 0.138 0.695 0.780
Gemini-2.0-Flash
Vanilla 0.045 0.145 0.831 0.223 0.253 0.577 0.654
Chain-of-Thought 0.042 0.118 0.837 0.232 0.255 0.558 0.663
Socratic Prompting 0.041 0.118 0.809 0.232 0.159 0.592 0.661
SubQ Decomposition 0.053 0.109 0.773 0.241 0.249 0.562 0.688
Claude-3.5-Sonnet
Vanilla 0.053 0.127 0.962 0.241 0.146 0.485 0.709
Chain-of-Thought 0.050 0.118 1.012 0.268 0.150 0.538 0.735
Socratic Prompting 0.044 0.118 0.998 0.232 0.145 0.508 0.748
SubQ Decomposition 0.050 0.136 0.990 0.259 0.149 0.485 0.741
Table A2: Evaluating top LLMs on T-FIX. We report the average performance of the LLM across all examples in the dataset. We report accuracy for classification tasks, and MSE for regression tasks – a () indicates that the score reported is MSE. Baseline implementations are described in Section 6.
Domain $\mathcal{N}$ generated claims $\mathcal{N}$ aligned claims Claim Decomposition Accuracy Relevance Filtering Accuracy Expert Alignment Accuracy Cohen’s $\kappa$
Cosmology
Mass Maps 66 48 0.900 0.826 0.979 0.4059
Supernova 74 62 0.950 0.892 0.903 0.4946
Psychology
Politeness 72 58 0.950 0.931 0.914 0.6604
Emotion 70 44 1.000 0.929 0.943 0.6233
Medicine
Cholecystectomy 134 92 1.000 0.851 0.902 0.4396
Cardiac 66 52 0.900 0.841 0.962 0.4845
Sepsis 108 66 0.900 0.852 0.894 0.3500
Table A3: Pipeline validation by domain. We report the mean accuracy for each stage of the pipeline and annotator agreement – Cohen’s $\kappa$.
Domain Claim Score (Category) Reasoning
Cosmology
Mass Maps [Good] The prominence of red and yellow suggests a universe with significant matter fluctuations. 0.9 (Density Contrast Extremes) Aligns well with the Density Contrast Extremes category, describing pronounced contrasts between dense and void regions, signaling high sigma_8.
[Bad] The mix of colors, with significant gray areas but noticeable reds and yellows, suggests a moderate Omega_m. 0.3 (Connectivity of the Cosmic Web) Discusses both underdense and overdense regions, but doesn’t specifically discuss connectivity or the degree of fragmentation or interconnection of the network.
Supernova [Good] A prominent peak followed by a gradual decline in flux is characteristic of a type Ia supernova light curve. 1.0 (Rise–decline rates) Describes a classic feature of type Ia supernovae, perfectly aligning with expert criteria on rise-and-decline rates.
[Bad] The variability does not display a clear periodicity. 0.1 (Periodic light curves) Contradicts key characteristics of periodic light curves; highlights absence of periodic behavior.
Psychology
Politeness [Good] The use of the phrase “seems defective” introduces uncertainty and avoids definitiveness. 0.9 (hedging & tentative language) The phrase utilizes tentative language and is a clear example of hedging to reduce the assertive strength of a statement.
[Bad] The utterance is a straightforward description of information from a biology textbook. 0.2 (First-Person Subjectivity Markers) Weakly aligns as it describes objective reporting without the personal tone central to first-person subjectivity.
Emotion [Good] This choice of description is likely intended to evoke a reaction of fear or caution. 0.9 (Threat/Worry Language) The claim centers around evoking fear or caution, which directly maps to this category.
[Bad] The text conveys an objective statement. 0.0 (Valence) The claim highlights an absence of emotional content, which does not align with the Valence category or any other expert emotion categories.
Medicine
Cholecys- [Good] The fat and fibrous tissue overlying Calot’s triangle has been fully excised, exposing only two tubular structures. High (Complete Triangle Clearance) Precisely describes complete clearance of Calot’s triangle, perfectly matching expert criteria.
tectomy [Bad] The cystic plate is not visible due to dense adhesions, making the gallbladder-liver plane indistinct. Low (Cystic Plate Visibility) Describes failure to visualize the cystic plate, opposite of the criterion, leading to low alignment.
Cardiac [Good] The irregularity in the ECG could indicate a dangerous arrhythmia, such as ventricular tachycardia or fibrillation. 0.9 (Ventricular Tachyarrhythmias) Directly references hallmark arrhythmias like ventricular tachycardia/fibrillation, key indicators in the category.
[Bad] A skin lesion of the scalp is a condition not directly related to cardiac function. 0.2 (Critical Illness – Sepsis/Shock) Potential weak connection if interpreted as infection, but lacks explicit signs of sepsis/shock.
Sepsis [Good] Fever and high heart rate are potential signs of sepsis. 1.0 (SIRS Positivity) References two SIRS criteria; strong and direct alignment with early sepsis identification guidelines.
[Bad] The patient’s lab results show an increased platelet count. 0.2 (SOFA Score Increase) SOFA score focuses on low platelet counts; increased count contradicts the criterion.
Table A4: Expert-aligned claims (good and bad) across all T-FIX domains, with corresponding alignment scores and provided reasoning.
Appendix A Extending T-FIX to a New Domain
Though T-FIX covers a wide range of knowledge-intensive settings, it can easily be extended to additional domains.
A key contribution of the T-FIX benchmark is the framework: we create a pipeline to score any free-form text explanation for expert alignment given a set of expert criteria. Additionally, we iterate extensively on all our prompt templates to ensure all T-FIX users need to do is input their task-specific details and perform no additional prompt engineering for good results.
To add a new domain to T-FIX, we advise you to follow these steps:
  1. 1.
    Generate criteria: Use the deep research prompt template shown in Figure˜A4 to generate a list of expert alignment criteria for your domain. Optionally, have a domain expert vet the generated criteria.
  2. 2.
    Modify prompts: Modify the prompt templates outlined in Figure˜A1, Figure˜A2, and Figure˜A3 with your task description, few-shot examples, and generated expert criteria.
  3. 3.
    Run T-FIX: Plug in your prompts for each stage of the pipeline and run T-FIX on your dataset!
We encourage you to contact the authors of this work if you need additional assistance setting up your custom domain.
Appendix B Prompts for T-FIX Pipeline
We show the prompts for Stage 1, 2, and 3 in Figure˜A1, Figure˜A2, and Figure˜A3, respectively. These prompts show a high-level template that was used by all domains. In practice, authors iterated multiple times on each domain’s prompts, experimenting with the instruction wording and few-shot examples that yielded the best possible results.
Appendix C Ablation for Evaluation Models
To assess the robustness of our evaluation pipeline, we conduct ablations using different models as evaluators: closed-sourced GPT-4o (openai2024gpt4ocard), Gemini-2.0-Flash (google2024gemini2flash), and an open-source model Qwen2.5-VL-7B-Instruct (Qwen2.5-VL). We evaluate vanilla prompts on three datasets, each representing a different domain: Mass Maps (cosmology), Emotion (psychology), and Cholecystectomy (medicine).
Table˜A1 summarizes the results:
  • Mass Maps (cosmology): All three evaluators unanimously rank Claude-3.5-Sonnet as the best model and o1 as the second best.
  • Cholecystectomy (medicine): GPT-4o and Qwen2.5-VL-7B-Instruct both rank o1 as the best and Qwen2.5-VL-7B-Instruct as the second best. Gemini-2.0-Flash is more conservative, labeling all claims as irrelevant and assigning all models a score of 0.
  • Emotion (psychology): All evaluators assign similar scores across generation models, leading to minimal differences in rankings.
Overall, these results indicate that our evaluation pipeline is robust to the choice of evaluation model.
Prompt
Figure A1: Prompt Template for Stage 1: Atomic Claim Extraction
Prompt
Figure A2: Prompt Template for Stage 2: Relevancy Filtering
Prompt
Figure A3: Prompt Template for Stage 3: Alignment Scoring
Prompt
Figure A4: Deep Research Prompt Template.
Prompt
Figure A5: Baseline Prompting Strategies.
Appendix D T-FIX Datasets: Additional Details
Prompt
Figure A6: MassMaps Explanation Prompt
D.1 Mass Maps
Task.
[…]
Data Selection & Preprocessing.
We randomly sampled 100 examples from the MassMaps test set. To ensure compatibility with LLMs like GPT-4o, which operate on a 32×32 patch size, we upsampled each image by a factor of 11 to preserve spatial detail and avoid patch-level compression. Instead of raw pixel values, we applied a colormap based on expert-defined intensity thresholds used to identify key cosmological features such as voids and clusters. Pixel intensities were scaled by standard deviations to emphasize meaningful variation. We found that larger, visually enhanced inputs reduced refusal rates from LLMs and encouraged more consistent responses.
Explanation Prompt.
Figure A6 shows the prompt used to generate LLM explanations for predicting $\Omega_{m}$ and $\sigma_{8}$. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of how pixel values are mapped to colors, as well as the valid ranges for $\Omega_{m}$ and $\sigma_{8}$. Without this range, models tend to default to common values (e.g., 0.3 for $\Omega_{m}$, 0.8 for $\sigma_{8}$), reducing response variability.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Lensing Peak (Cluster) Abundance: High peak count $\rightarrow$ higher $\sigma_{8}$; clumpy halos more common.
  2. 2.
    Void Size and Frequency: Large, frequent voids $\rightarrow$ lower $\Omega_{m}$; less overall matter.
  3. 3.
    Filament Thickness and Sharpness: Thick, sharp filaments track higher $\sigma_{8}$; thin indicates lower.
  4. 4.
    Fine-Scale Clumpiness: Fine graininess signifies high $\sigma_{8}$; smooth map implies lower.
  5. 5.
    Connectivity of the Cosmic Web: Interconnected web suggests higher $\Omega_{m}$; isolated clumps imply lower.
  6. 6.
    Density Contrast Extremes: Strong density contrast denotes high $\sigma_{8}$; muted contrast lower.
Prompt
Figure A7: Supernova Explanation Prompt
D.2 Supernova
Task.
The objective is to classify astrophysical objects using time-series data comprising observation times (Modified Julian Dates), wavelengths (filters), flux values, and corresponding flux uncertainties. We use data from the PLAsTiCC challenge (theplasticcteam2018photometric), where the model must predict one of 14 astrophysical classes.
Data Selection & Preprocessing.
We sampled 100 examples across the Supernova train, validation, and test sets, aiming for 7–8 instances per class to mitigate class imbalance. For rare classes with only one test set instance, we included all available examples from the validation and test sets, supplementing with training samples to meet the target count. For LLM input, we converted each raw time series into a multivariate time-series plot: time is on the x-axis, flux on the y-axis, error bars denote flux uncertainty, and point colors indicate different wavelengths.
Explanation Prompt.
Figure A7 shows the prompt used to generate explanations for classifying astronomical objects. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of the input plot as a multivariate time series and provides the full list of possible class labels to guide the model’s predictions.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Contiguous non-zero flux: Contiguous non‑zero flux segments confirm genuine astrophysical activity and define the time windows from which transient features should be extracted.
  2. 2.
    Rise–decline rates: Characteristic rise‑and‑decline rates—such as the fast‑rise/slow‑fade morphology of many supernovae—encode energy‑release physics and serve as strong class discriminators.
  3. 3.
    Photometric amplitude: Peak‑to‑trough photometric amplitude separates high‑energy explosive events (multi‑magnitude outbursts) from low‑amplitude periodic or stochastic variables.
  4. 4.
    Event duration: Total event duration, measured from first detection to return to baseline, distinguishes short‑lived kilonovae and superluminous SNe from longer plateau or AGN variability phases.
  5. 5.
    Periodic light curves: Periodic light curves with stable periods and distinctive Fourier amplitude‑ and phase‑ratios flag pulsators and eclipsing binaries rather than one‑off transients.
  6. 6.
    Secondary maxima: Filter‑specific secondary maxima or shoulders in red/near‑IR bands—prominent in SNeIa—are morphological features absent in most core‑collapse SNe.
  7. 7.
    Monotonic flux trends: Locally smooth, monotonic flux trends across one or multiple bands (plateaus, linear decays) capture physical evolution stages and help distinguish SNII‑P, SNII‑L, and related classes.
D.3 Politeness
Task.
Understanding how linguistic styles, like politeness, vary across cultures is necessary for building better communication, translation, and conversation-focused systems. holmes2012politeness; havaldar2023multilingual. Today’s LLMs exhibit large amounts of cultural bias havaldar2024building, and understanding nuances in cultural differences can help encourage cultural adaptation in models. We use the holistic politeness dataset from havaldar2025comparingstyleslanguagescrosscultural, which consists of conversational utterances between editors from Wikipedia talk pages, annotated by native speakers from four distinct cultures.
Data Selection & Preprocessing.
We sample 100 examples from the data, balanced equally across classes (rude, slightly rude, neutral, slightly polite, polite) and languages (English, Spanish, Japanese, Chinese).
Prompt
Figure A8: Politeness Explanation Prompt
Explanation Prompt.
We show the prompt in Figure˜A8. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Honorifics and Formal Address: The presence of respectful or formal address forms (e.g., “sir,” “usted,”) signals politeness by expressing deference to the hearer’s status or social distance.
  2. 2.
    Courteous Politeness Markers: Words such as “please,” “kindly,” or their multilingual variants soften requests and reflect courteous intent.
  3. 3.
    Gratitude Expressions: Use of expressions like “thank you,” “thanks,” or “I appreciate it” signals recognition of the other’s contribution and positive face.
  4. 4.
    Apologies and Acknowledgment of Fault: Phrases such as “sorry” or “I apologize” express humility and repair social breaches, marking a clear politeness strategy.
  5. 5.
    Indirect and Modal Requests: Requests using modal verbs (“could you,” “would you”) or softening cues like “by the way” reduce imposition and signal respect for the hearer’s autonomy.
  6. 6.
    Hedging and Tentative Language: Words like “I think,” “maybe,” or “usually” lower assertion strength and make statements more negotiable, reflecting interpersonal sensitivity.
  7. 7.
    Inclusive Pronouns and Group-Oriented Phrasing: Use of “we,” “our,” or “together” expresses solidarity and reduces hierarchical distance in requests or critiques.
  8. 8.
    Greeting and Interaction Initiation: Opening with a salutation (“hi,” “hello”) creates a cooperative tone and frames the conversation positively.
  9. 9.
    Compliments and Praise: Positive evaluations (“great,” “awesome,” “neat”) attend to the hearer’s positive face and foster a friendly environment.
  10. 10.
    Softened Disagreement or Face-Saving Critique: When disagreeing, the use of softeners, partial agreements, or concern for clarity preserves the hearer’s dignity.
  11. 11.
    Urgency or Immediacy of Language: Utterances emphasizing emergency or speed (“asap,” “immediately”) can heighten perceived imposition and reduce politeness if not softened.
  12. 12.
    Avoidance of Profanity or Negative Emotion: The presence of strong negative words or swearing is a key indicator of rudeness and face threat.
  13. 13.
    Bluntness and Direct Commands: Requests lacking modal verbs or mitigation (“Do this”) are perceived as less polite due to their imperative structure.
  14. 14.
    Empathy or Emotional Support: Recognizing the hearer’s emotional context or challenges is a politeness strategy of concern and goodwill.
  15. 15.
    First-Person Subjectivity Markers: Statements that begin with “I think,” “I feel,” or “In my view” convey humility and subjectivity, reducing imposition.
  16. 16.
    Second Person Responsibility or Engagement: Sentences starting with “you” or directly addressing the hearer can either signal engagement or come across as accusatory, depending on context and tone.
  17. 17.
    Questions as Indirect Strategies: Questions (“what do you think?” or “could you clarify?”) reduce imposition by inviting rather than demanding input.
  18. 18.
    Discourse Management with Markers: Use of discourse markers like “so,” “then,” “but” organizes conversation flow and may help manage face needs in conflict or negotiation.
  19. 19.
    Ingroup Language and Informality: Use of group-identifying slang or casual expressions (“mate,” “dude,” “bro”) may foster solidarity or seem disrespectful, depending on relational norms.
D.4 Emotion
Task.
Understanding and classifying emotion is important for tasks like therapy, mental health diagnoses, etc. denzin1984understanding. Emotion is often expressed implicitly, and understanding such cues can also aid in building LLM systems that handle implied language understanding well havaldar2025entailed. We use the GoEmotions dataset from demszky2020goemotions, consisting of Reddit comments that have been human-annotated for one of 27 emotions (or neutral, if no emotion is present).
Data Selection & Preprocessing.
We sample 100 examples from the data, balanced equally across 28 emotion classes, including neutral. We additionally ensure the comment is over 20 characters, to remove noisy data points and ensure each comment contains enough information for the LLM to make an accurate classification.
Prompt
Figure A9: Emotion Explanation Prompt
Explanation Prompt.
We show the prompt in Figure˜A9. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Valence: Decide if the overall tone is pleasant or unpleasant; positive tones suggest joy or admiration, negative tones suggest sadness or anger.
  2. 2.
    Arousal: Gauge how energized the wording is—calm phrasing implies low arousal emotions, intense phrasing implies high arousal emotions.
  3. 3.
    Emotion Words & Emojis: Look for direct emotion terms or emoticons that explicitly name the feeling.
  4. 4.
    Expressive Punctuation: Multiple exclamation marks, ALL‑CAPS, or stretched spellings signal higher emotional intensity.
  5. 5.
    Humor/Laughter Markers: Tokens like “haha,” “lol,” or laughing emojis reliably indicate amusement.
  6. 6.
    Confusion Phrases: Statements such as “I don’t get it” clearly mark confusion.
  7. 7.
    Curiosity Questions: Genuine information‑seeking phrases (“I wonder…”, “why is…?”) point to curiosity.
  8. 8.
    Surprise Exclamations: Reactions of astonishment (“No way!”, “I can’t believe it!”) denote surprise.
  9. 9.
    Threat/Worry Language: References to danger or fear (“I’m scared,” “terrifying”) signal fear or nervousness.
  10. 10.
    Loss or Let‑Down Words: Mentions of loss or disappointment cue sadness, disappointment, or grief.
  11. 11.
    Other‑Blame Statements: Assigning fault to someone else for a bad outcome suggests anger or disapproval.
  12. 12.
    Self‑Blame & Apologies: Admitting fault and saying “I’m sorry” marks remorse.
  13. 13.
    Aversion Terms: Words like “gross,” “nasty,” or “disgusting” point to disgust.
  14. 14.
    Praise & Compliments: Positive evaluations of someone’s actions show admiration or approval.
  15. 15.
    Gratitude Expressions: Phrases such as “thanks” or “much appreciated” indicate gratitude.
  16. 16.
    Affection & Care Words: Loving or nurturing language (“love this,” “sending hugs”) signals love or caring.
  17. 17.
    Self‑Credit Statements: Boasting about one’s own success (“I nailed it”) signals pride.
  18. 18.
    Relief Indicators: Release phrases like “phew,” “finally over,” or “what a relief” mark relief after stress ends.
D.5 Laparoscopic Cholecystectomy Surgery.
Task.
The task is to identify the safe and unsafe regions for incision. We used the open-source subset of data from (madani2022artificial), which consists of surgeon-annotated images taken from video frames from the M2CAI16 workflow challenge (stauder2016tum) and Cholec80 (twinanda2016endonet) datasets. This consists of 1015 surgeon-annotated images.
Data Selection & Preprocessing.
We selected the first 100 items from the test set where the safe and unsafe regions were of nontrivial area. Each item has three components: an image of dimensions 640 pixels wide by 360 pixels high, a binary mask of the safe regions of the same dimensions, and a binary mask of the unsafe regions of the same dimensions.
To convert the task into a form easily solvable by the available APIs, our objective was to have the LLM output a small list of numbers that identify the safe and unsafe regions. This is achieved by using square grids of size 40 to discretize each of the safe and unsafe masks, separating them into $144=(640/40)\times(360/40)$ disjoint regions. One can then use an integer inclusively ranging from 0 to $143$ to uniquely identify these patches. The LLM was to then output two lists with numbers from this range: a “safe list” that denotes its prediction of the safe region, and an “unsafe list” predicting the unsafe region.
Explanation Prompt.
We show the prompt in Figure˜A10. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Calot’s triangle cleared - Hepatocystic triangle must be fully cleared of fat/fibrosis so that its boundaries are unmistakable.
  2. 2.
    Cystic plate exposed - The lower third of the gallbladder must be dissected off the liver to reveal the shiny cystic plate and ensure the correct dissection plane.
  3. 3.
    Only two structures visible - Only the cystic duct and cystic artery should be seen entering the gallbladder before any clipping or cutting.
  4. 4.
    Above the R4U line - Dissection must remain cephalad to an imaginary line from Rouviere’s sulcus to liver segment IV to avoid the common bile duct.
  5. 5.
    Safe distance from common bile duct - There should be sufficient distance between the common bile duct and the gallbladder wall to ensure safe dissection.
  6. 6.
    Infundibulum start point - Dissection should begin at the gallbladder infundibulum-cystic duct junction to stay in safe tissue planes.
  7. 7.
    Subserosal plane stay - When separating the gallbladder from the liver, stay in the avascular subserosal cleavage plane under the serosal fat layer.
  8. 8.
    Cystic lymph node guide - Identify the cystic lymph node and clip the artery on the gallbladder side of the node to avoid injuring the hepatic artery.
  9. 9.
    No division without ID - Never divide any duct or vessel until it is unequivocally identified as the cystic structure entering the gallbladder.
  10. 10.
    Inflammation bailout - If dense scarring or distorted anatomy obscures Calot’s triangle, convert to a subtotal "fundus-first" approach rather than blind cutting.
  11. 11.
    Aberrant artery caution - Preserve any large or tortuous artery (e.g., a Moynihan’s hump) that might be mistaken for the cystic artery.
Prompt
Figure A10: Laparoscopic Cholecystectomy Explanation Prompt. A list of 10 few-shot examples is then appended to the same API call. Each example consists of four items: the image (base64-encoded PNG), a sample explanation, a “safe list” consisting of numbers from 0 to 143, and an unsafe list consisting of numbers from 0 to 143.
Prompt
Figure A11: Cardiac Explanation Prompt
D.6 Cardiac Arrest
Task.
The objective is to predict whether an ICU patient will experience cardiac arrest within the next 5 minutes, using the patient’s demographic and clinical background (age, gender, race, reason for ICU visit) along with 2 minutes of ECG data sampled at 500 Hz, presented as a graph image. This framing aligns with cardiology literature, which suggests that short ECG windows (30 seconds to a few minutes) are sufficient for reliable prediction nussinovitch2011reliability. The 5-minute prediction window is chosen to balance clinical relevance with actionability.
Data Selection & Preprocessing.
We use ECG and visit data from the open-source Multimodal Clinical Monitoring in the Emergency Department (MC-MED) Dataset kansal2025mcmed. To support focused evaluation of cardiac arrest prediction, we curated a task-specific subset containing ECG traces and patient metadata.
The data curation pipeline proceeded as follows. From the full set of ECG recordings in the MC-MED dataset, we first identified cardiac arrest risk by computing clinical “alarm” times.
Prior work shows that vital sign abnormalities are predictive of outcomes Candel2022-bo; chen2023multimodal. We defined an alarm at any timestamp where three or more of the following vital signs were outside normal range within a two-minute window—a condition known clinically as decompensation:
  • Heart rate (HR): < 40 or > 130 bpm
  • Respiratory rate (RR): < 8 or > 30 breaths/min
  • Oxygen saturation (SpO2): < 90%
  • Mean arterial pressure (MAP): < 65 or > 120 mmHg
Each example was labeled ’Yes’ if an alarm was present, and ’No’ otherwise. For positive cases, we sampled a random cutoff time 1–300 seconds before the alarm and extracted the preceding 2 minutes of ECG data. For negative cases, we used the first 2 minutes of ECG data. We also added patient metadata—age, gender, race, and ICU admission reason—using information from the MC-MED visit records. To ensure diversity, each example came from a unique patient; for positives, we only used the visit containing the alarm.
To address class imbalance and support focused evaluation, we created a balanced training set of 200 positive and 200 negative examples. The validation and test sets each contain 50 examples.
Explanation Prompt.
Figure A11 shows the prompt used to generate explanations for predicting whether an ICU patient will experience cardiac arrest within 5 minutes, based on 2 minutes of ECG data along with age, gender, race, and ICU admission reason. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The ECG is provided as a graph image of p-signal values sampled at 500 Hz over a 2-minute window, with labeled axes. While we considered supplying the raw signal as text, the input token limits of current LLMs made this infeasible.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Ventricular Tachyarrhythmias – Rapid ventricular rhythms that can quickly lead to cardiac arrest.
  2. 2.
    Ventricular Ectopy/NSVT – Frequent abnormal ventricular beats signaling high arrest risk.
  3. 3.
    Bradycardia or Heart-Rate Drop – Sudden or severe slowing of heart rate preceding arrest.
  4. 4.
    Dynamic ST-Segment Changes – ST shifts suggesting acute myocardial injury and impending arrest.
  5. 5.
    Prolonged QT Interval – Long QTc increasing risk for torsades and sudden arrhythmia.
  6. 6.
    Severe Hyperkalemia Signs – ECG changes from high potassium predicting arrest, especially among patients on dialysis / end stage renal disease.
  7. 7.
    Advanced Age – Older age strongly correlates with higher arrest likelihood.
  8. 8.
    Male Sex – Males have a higher overall risk of cardiac arrest.
  9. 9.
    Underlying Cardiac Disease – Preexisting heart disease increases arrest susceptibility.
  10. 10.
    Critical Illness (Sepsis/Shock) – Severe infections or shock states elevate arrest risk through systemic instability.
Prompt
Figure A12: Sepsis Explanation Prompt
D.7 Sepsis
Task.
The goal is to predict whether an emergency department (ED) patient is at high risk of developing sepsis within 12 hours, using Electronic Health Record (EHR) data collected during the first 2 hours of their visit. Each input is a time series of records containing a timestamp, the name of a physiological measurement or medication, and its value.
Data Selection & Preprocessing.
We used data from the publicly available MC-MED dataset kansal2025mcmed and curated a task-specific subset for sepsis prediction.
To label a patient as high risk for sepsis, we followed standard clinical definitions requiring three conditions: (1) evidence of infection, indicated by either a blood culture being drawn or at least two hours of antibiotic administration; (2) signs of organ dysfunction, defined by a SOFA score $\geq$2 within 48 hours of suspected infection, based on abnormalities in respiratory, coagulation, liver, cardiovascular, neurological, or renal function; and (3) presence of fever, with a recorded temperature $\geq$38.0°C (100.4°F). Patients meeting all three criteria were labeled as high risk. Labels were validated with a Sepsis clinician.
Due to class imbalance (1̃0% positive), we created a balanced evaluation set of 100 samples (50 positive, 50 negative) drawn from the validation and test splits.
Explanation Prompt.
Figure A12 shows the prompt used to generate LLM explanations for sepsis risk prediction. We substitute [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of the EHR input format: each time-series record consists of a timestamp, a measurement or medication name, and its value.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
  1. 1.
    Elderly Susceptibility (Age $\geq$65 years): Advanced age ($\geq$65 years) markedly increases susceptibility to rapid sepsis progression and higher mortality after infection.
  2. 2.
[…]
  3. 3.
    High qSOFA Score ($\geq$2): A qSOFA score $\geq$2 (respiratory rate $\geq$22/min, systolic BP $\leq$100 mmHg, or altered mentation) flags high risk of sepsis-related organ dysfunction and mortality.
  4. 4.
    Elevated NEWS Score ($\geq$5 points): A National Early Warning Score (NEWS) of $\geq$5–7 derived from deranged vitals predicts imminent clinical deterioration compatible with sepsis.
  5. 5.
    Elevated Serum Lactate ($\geq$2 mmol/L): Serum lactate $\geq$2 mmol/L within the first 2 hours signals tissue hypoperfusion and markedly elevates sepsis mortality risk.
  6. 6.
    Elevated Shock Index ($\geq$1.0): Shock index (heart rate $\div$ systolic BP) $\geq$1.0—or a rise $\geq$0.3 from baseline—denotes haemodynamic instability and a high probability of severe sepsis.
  7. 7.
    Sepsis-Associated Hypotension (SBP <<90 mmHg or MAP <<70 mmHg, or $\geq$40 mmHg drop): Sepsis-associated hypotension, defined as SBP <<90 mmHg, MAP <<70 mmHg, or a $\geq$40 mmHg drop from baseline, indicates progression toward septic shock.
  8. 8.
    SOFA Score Increase ($\geq$2 points): An increase of $\geq$2 points in any SOFA component—e.g., PaO2/FiO2 <<300, platelets $0×\times1$09/L, bilirubin >>2 mg/dL, creatinine >>2 mg/dL, or GCS <<12—confirms new organ dysfunction and high sepsis risk.
  9. 9.
    Early Antibiotic/Culture Orders (within 2 hours): Administration of broad-spectrum antibiotics or drawing of blood cultures within the first 2 hours signifies clinician suspicion of serious infection and should anchor sepsis risk assessment.
Appendix E Expert Validation
E.1 Mass Maps
Experts largely agreed with the model’s alignment on simple, descriptive claims that focused on observable visual patterns rather than interpretive reasoning. For instance, statements such as “The map displays a noticeable amount of blue and gray with significant red clustering” or mentions of “yellow spots” as dense regions were seen as well-aligned because they accurately reflected direct features in the image without overinterpreting them. The model was particularly consistent in correctly associating visible peaks and clusters with the relevant “Lensing Peak (Cluster) Abundance” category, showing strong performance on basic lensing-related observations.
However, disagreements arose for more interpretive claims where the model either misclassified the category or underestimated alignment. For example, experts judged “The significant red clustering indicates areas with some dense structures” as completely aligned, while the model labeled it partial. Similarly, when the model associated “yellow spots” with fine-scale clumpiness instead of large cluster convergence, experts noted categorical confusion.
At a higher level, experts appreciated the model’s handling of concrete, color- or peak-based features but found its phrasing vague when dealing with broader cosmological interpretations. They preferred precise terminology about structure type and scale, emphasizing that generic descriptions such as “noticeable fluctuations” failed to convey sufficient scientific specificity.
E.2 Supernova
Experts agreed with alignment scores for claims describing basic observable behaviors, such as “The light curve is followed by a gradual decline” or “The multi-wavelength observation shows a subsequent decline across all wavelengths,” which correctly corresponded to the “Monotonic Flux Trend” category. These statements were straightforward and descriptive, and the system’s partial alignment assessments matched expert expectations.
Disagreement arose for claims that overinterpreted observational patterns. For example, “The decline across all wavelengths corroborates the classification as a Type Ia supernova” was given full alignment by the model, but experts argued it should be partial, as such behavior is characteristic of transient events generally, not specific to Type Ia supernovae.
Experts praised the model’s ability to identify key distinguishing features, like “The light curve shows a rapid rise to a peak,” which effectively differentiates supernovae from other celestial phenomena. However, they were less satisfied with claims that only ruled out other classes (e.g., RR Lyrae, AGN, Mira) without affirmatively justifying the target classification. They stressed that negative reasoning — describing what an instance is not — does not equate to correctly identifying what it is.
E.3 Politeness
Experts generally agreed with the model’s alignment on explicit and lexical politeness cues. For example, both the model and experts assigned full alignment to claims like “The speaker expresses appreciation by saying ‘thank you,’” and “The utterance softens the request through the word ‘please.’” These cases show that the model reliably recognizes direct politeness indicators that have stable, context-independent meanings.
Disagreement arose for subtler forms of politeness that rely on social context or pragmatic interpretation. Experts noted that the model often overestimated alignment for hedges and modal constructions, such as I think, maybe, or could you, which did not always convey genuine politeness signal. For instance, the model rated “The use of ‘we’ conveys solidarity” as fully aligned, while experts marked it partial because the “we” referred to an institution rather than shared identity.
At a higher level, experts appreciated the model’s consistency and strong performance on overt politeness expressions but emphasized its limited pragmatic sensitivity. Overall, experts found the LLM trustworthy for identifying surface politeness cues but less dependable when interpreting indirect, contextual, or culturally grounded politeness strategies.
E.4 Emotion
Experts generally agreed with the model’s alignment for surface-level emotional cues. Claims referencing Emotion Words, Emojis, Expressive Punctuation, Humor Markers were accurately scored by both the model and experts. The model also showed strong reliability in labeling claims related to Gratitude and Praise. These results suggest that the model can be trusted to judge alignment accurately when the emotion is lexically or visually explicit.
Disagreement emerged in more nuanced emotional expressions where context determined the emotional interpretation. Experts noted that the model often overestimated alignment for Valence and Arousal categories, particularly when phrasing was neutral or ambiguous. Claims involving subtle or mixed emotions, such as those involving Relief or Affection, were rarely labeled partial, as the model treated any positive cue as a strong emotional expression.
At a broader level, experts were happy with the model’s precision in labeling but noted limited sensitivity to gradation and context. The system performed best when emotions were unambiguous and directly expressed, but faltered when affect was implied, ironic, or intertwined with multiple sentiments.
E.5 Cardiac
Experts found most alignment scores accurate and consistent with their own judgments, highlighting the model’s general reliability in this domain. They appreciated the structural clarity and systematic nature of the pipeline, noting that alignment scores often captured correct relationships between claims and expert categories. Nevertheless, they pointed out that while scores were often correct, the model’s reasoning behind them was sometimes incomplete or slightly off, revealing a disconnect between the correctness of the label and the explanatory rationale.
At a broader level, experts valued the coherence of the approach and the alignment consistency across claims. However, they criticized redundancy and excessive complexity in some claims. For example, two statements about ECG rhythm regularity were nearly identical, suggesting overgeneration. They also noted that some claims were too dense and should be broken into smaller, more specific parts. Additional concerns included inaccuracies in certain category definitions, such as labeling 30 years old as “advanced age”, and insufficient background context for some explanations.
Overall, they viewed the system as strong in structure but in need of refinement in content granularity and contextual accuracy.
E.6 Sepsis
Experts agreed with most alignment scores, particularly those labeled partial, since the claims captured correct but incomplete aspects of the clinical reasoning process. For instance, “One risk factor for sepsis is advanced age” was appropriately labeled as partial under “Elderly Susceptibility,” as it was factually true but lacked the specific clinical threshold (age $\geq$ 65). Similarly, claims about single vital sign abnormalities were correctly labeled as partial since sepsis criteria like SIRS require multiple indicators.
However, experts disagreed when the model evaluated isolated claims that should have been considered in combination. Statements like “A high triage temperature indicates fever” and “An elevated respiratory rate is another risk factor” were accurate but insufficient alone to establish “SIRS Positivity,” leading to misalignment between formal criteria and contextual reasoning. Additionally, experts highlighted that the LLM would generate explanations that involved risk factors that are not the same as diagnostic criteria for sepsis — while risk factors increase the likelihood of sepsis, there is no causal relationship.
At a higher level, experts appreciated the clarity of claims tied to objective measures such as qSOFA scores and age but highlighted a persistent tension between rule-based alignment and real-world clinical reasoning. They emphasized that diagnostic logic in practice is more flexible than rigid scoring systems, exposing a conceptual gap that the model did not fully capture.

### v2

@@
 Abstract
-As LLMs are deployed in knowledge‑intensive settings (e.g., surgery, astronomy, therapy), users expect not just answers, but also meaningful explanations for those answers.
-In these settings, users are often domain experts (e.g., doctors, astrophysicists, psychologists) who require explanations that reflect expert-level reasoning.
-However, current evaluation schemes primarily emphasize plausibility or internal faithfulness of the explanation, which fail to capture whether the content of the explanation truly aligns with expert intuition.
-We formalize expert alignment as a criterion for evaluating explanations with T‑FIX, a benchmark spanning seven knowledge-intensive domains.
-In collaboration with domain experts, we develop novel metrics to measure the alignment of LLM explanations with expert judgment.
-[^1] https://github.com/BrachioLab/FIX-2
-T-FIX: Text-Based Explanations with Features Interpretable to eXperts
-[…]
+As LLMs are deployed in knowledge-intensive settings (e.g., surgery, astronomy, therapy), users are often domain experts who expect not just answers, but explanations that mirror professional reasoning.
+However, most automatic evaluations of explanations prioritize plausibility or faithfulness, rather than testing whether an LLM “thinks like an expert.” Existing approaches to evaluating professional reasoning rely heavily on per-example expert annotation, making such evaluations costly and difficult to scale.
+To address this gap, we introduce the T-FIX benchmark, spanning seven scientific tasks across three domains, to operationalize expert alignment as a desired attribute of LLM-generation explanations.
+Our framework enables automatic evaluation of expert alignment, generalizing to unseen explanations and eliminating the need for ongoing expert involvement.
+Code and data available at https://github.com/BrachioLab/FIX-2
+Machine Learning, ICML
 1 Introduction
-Figure 1: Current evaluations of LLM explanations typically consider two dimensions: plausibility — whether the reasoning is logically coherent and faithfulness — whether it reflects the model’s true decision process.
-We introduce a third, orthogonal dimension: expert alignment — whether the LLM reasons as a domain expert would.
-For instance, an LLM may correctly predict sepsis risk with a plausible and faithful explanation, yet because it relies on features clinicians rarely use, its expert alignment is low.
+Figure 1: Current automatic evaluations of LLM explanations focus on plausibility — whether the reasoning is logically coherent and faithfulness — whether it reflects the model’s true decision process.
+We formalize a third, orthogonal dimension: expert alignment — whether the LLM reasons as a domain expert would.
+For instance, an LLM may correctly predict sepsis risk, yet because it relies on features clinicians rarely use, its expert alignment is low.
 Figure 2: An overview of the T-FIX construction process.
@@
 A high score suggests the explanation reflects reasoning aligned with domain experts (i.e., the LLM “thinks like an expert”), while a low score indicates the explanation may rely on aspects that experts would deem irrelevant.
-LLMs are increasingly used for domain-specific tasks that require substantial knowledge from specialized fields — the adoption of LLMs is actively being explored in high-stakes environments like operating rooms, astronomical observatories, and therapy clinics pressman2024clinical; fouesneau2024rolelargelanguagemodels; stade2024large.
-For LLMs to be trustworthy in these critical applications, users require not only correct answers but also good explanations rudin2019stop; pedreschi2019meaningful.
-However, what constitutes a “good explanation” depends on the target audience ribera2019can; sokol2020one.
-As LLMs are used for specialized tasks like surgical assistance or supernova analysis, the primary users are often domain experts, like doctors or astrophysicists.
-Consequently, a “good explanation” in these specialized contexts must offer insights that are valuable and interpretable to these domain experts.
-Existing evaluations of LLM explanations focus primarily on two dimensions: (1) plausibility, which assesses whether an explanation appears intuitively reasonable to an average human observer; and (2) faithfulness, which measures whether the explanation accurately reflects the LLM’s internal reasoning process zhou2021evaluating; agarwal2024faithfulness; parcalabescu2023measuring.
-While these dimensions are necessary, they are insufficient for knowledge-intensive domains.
-Experts do not merely ask whether an explanation seems plausible or mirrors the model’s internal reasoning; they require that the reasoning itself aligns with domain-validated principles wang2021explanations.
-In such settings, the key question is whether LLMs consider aspects of the input that experts themselves deem scientifically critical.
-This limitation is illustrated in Figure 1: an LLM may generate a correct answer with a plausible and faithful explanation, yet still rely on features irrelevant to the expert’s reasoning framework.
-Such reasoning, though logically coherent, fails to meet expert standards.
-To address this, we introduce a third evaluation dimension: expert alignment.
-Expert alignment assesses whether an explanation emphasizes the criteria a domain expert would use when making the same prediction.
-It is orthogonal to plausibility and faithfulness: while plausibility evaluates logical coherence and faithfulness evaluates causal accuracy, expert alignment tests for epistemic validity — that the explanation is grounded in established scientific knowledge.
-While prior work has examined expert reasoning via structured feature attribution jin2024fix, such methods are largely confined to classification settings.
-In contrast, modern LLMs produce free-form textual explanations not directly linked to explicit feature groups.
-To our knowledge, no existing benchmark evaluates the alignment of such free-form explanations.
-To fill this gap, we introduce the T-FIX benchmark: a suite of datasets spanning seven domains with an accompanying evaluation framework.
-Developed with domain experts, T-FIX quantifies the extent to which LLM explanations align with scientifically-grounded expert reasoning.
+LLMs are increasingly used for domain-specific tasks that require substantial knowledge from specialized fields — the adoption of LLMs is actively being explored in high-stakes environments like operating rooms, astronomical observatories, and therapy clinics (Pressman et al., 2024; Fouesneau et al., 2024; Stade et al., 2024).
+For LLMs to be trustworthy in these critical applications, users require not only correct answers but also good explanations (Rudin, 2019; Pedreschi et al., 2019).
+However, what constitutes a “good explanation” depends on the target audience (Ribera and Lapedriza, 2019; Sokol and Flach, 2020).
+In specialized tasks such as surgical assistance or supernova analysis, the primary users are domain experts (e.g., doctors or astrophysicists).
+“Thus, a good explanation” must surface the scientifically relevant criteria experts use in practice.
+This paper introduces T-FIX to measure this at scale: a suite of datasets and an expert-designed evaluation framework that automatically measures how much an explanation aligns with expert intuition and generalizes to unseen examples across scientific domains.
+Current automatic evaluations of LLM explanations typically focus on two dimensions: (1) plausibility, whether an explanation appears intuitively reasonable to an average observer; and (2) faithfulness, whether it reflects the LLM’s internal reasoning process (Zhou et al., 2021; Agarwal et al., 2024; Parcalabescu and Frank, 2023).
+While necessary, these dimensions are insufficient in knowledge-intensive domains, where experts require reasoning that aligns with domain-validated principles (Wang and Yin, 2021); the key question is whether LLMs attend to aspects of the input that experts deem scientifically critical.
+As Figure 1 illustrates, an LLM may produce a correct answer with a plausible and faithful explanation, yet still rely on features irrelevant to expert reasoning; such explanations are coherent but fail to meet expert standards.
+We argue for automatic evaluation of a third dimension, expert alignment, which measures whether an explanation emphasizes the criteria a domain expert would use, and tests epistemic validity — grounding in established scientific knowledge, rather than coherence or causal fidelity.
+Current methods for measuring expert alignment of LLM outputs across medicine, psychology, and astronomy are not automatic or scalable because they require substantial manual labor from the expert.
+For example, Kutbi et al.
+(2025) annotate only 18 sepsis-diagnosis explanations with 7 physicians, and Hyk et al.
+(2025) report four weeks of interviews with 11 astronomers.
+These studies are difficult to turn into benchmarks: the evaluations are limited in size and often qualitative, and therefore do not generalize beyond the specific annotated examples.
+Conversely, larger benchmarks for domain-specific LLM evaluation rarely assess expert alignment in a substantive way.
+In many cases, experts appear primarily in validation or annotation studies (e.g., Astro-QA (Li et al., 2025a), CounselBench (Li et al., 2025b), Mentalchat16k (Xu et al., 2025)).
+In medicine, newer efforts introduce per-example, domain-specific rubrics (e.g., Health-SCORE from Yang et al.
+(2026) and OpenAI’s HealthBench), but they share a key limitation: expert input is tied to the benchmark’s fixed examples and does not readily transfer to evaluating explanations on unseen cases.
+To fill this gap, we introduce the T-FIX benchmark: a suite of seven datasets spanning three scientific domains with an accompanying expert-designed evaluation framework.
+Developed with domain experts, T-FIX quantifies how well LLM explanations align with scientifically grounded expert reasoning.
 Our contributions are as follows:
-  • We introduce expert alignment as a desired attribute of LLM-generated explanations and create T-FIX, the first benchmark designed to evaluate this.
-  • We release a pipeline to evaluate how well any LLM “thinks like an expert,” designed to be easily extendable to new domains.
-  • We show that while LLMs perform better when reasoning over multiple expert criteria, they still struggle to align their explanations with expert intuition, revealing a key direction for future improvement.
-2 Expert Alignment Criteria
+  • We formalize expert alignment as a desired attribute of LLM-generated explanations and create T-FIX, the first benchmark designed to evaluate this automatically.
+  • We release an expert-designed framework to evaluate how well any LLM “thinks like an expert,” across seven unique scientific tasks.
+  • We show that while LLMs perform better when reasoning over multiple expert criteria, they do not explicitly rely on expert reasoning, revealing a key direction for future improvement.
+T-FIX & Scalability.
+With T-FIX, expert effort is front-loaded and reduced: experts help develop a single, robust evaluation framework, after which users can score expert alignment on unseen LLM explanations automatically, eliminating the need for future per-example annotations or additional expert time.
+2 Expert Criteria for Evaluating Explanations
 Figure 3: Our T-FIX pipeline.
 To evaluate an LLM-generated explanation, we first decompose it into atomic claims.
-Next, we filter out irrelevant claims, such as unsupported or speculative statements.
-Each remaining claim is then scored against the domain-specific expert alignment criteria: a score of “complete” indicates perfect overlap with at least one criterion, while “none” indicates no overlap.
-Filtered-out claims are automatically assigned a score of “none”.
-We compute the final expert-alignment score for the explanation by averaging across all claim scores.
+Next, we filter out irrelevant claims (e.g., unsupported or speculative statements) and group the remaining claims by their relation to each expert criterion.
+We then score each criterion-level group against its corresponding expert criterion as complete, partial, or none.
+Each claim inherits the highest score among the groups it belongs to, while filtered-out or ungrouped claims are assigned none.
+We compute the final expert alignment score by averaging these per-claim scores across the explanation.
 The development of the T-FIX benchmark was a highly collaborative and interdisciplinary process.
 For each of our seven domains (see Figure 4), our first step was to identify the expert criteria most relevant to making a prediction, detailed in the left of Figure 2.
-When answering knowledge-intensive questions like “Will this patient develop sepsis in the next 12 hours?” or “What kind of supernova produced these wavelengths?”, doctors and astrophysicists rely on domain-specific heuristics, prioritizing certain features over others based on training and experience.
-For instance, in sepsis classification, an experienced clinician would typically emphasize features like advanced age and hypotension, while assigning lower importance to signals like glucose levels or patient demeanor, which are less directly indicative of sepsis risk.
-Thus, an LLM that makes the correct prediction by attending to age and hypotension is more expert-aligned than one that arrives at the same answer by focusing on glucose and demeanor.
-We define the subset of features that experts prioritize when performing a task as expert alignment criteria.
-Step 1: Surveying the field.
-To seed our initial list of expert criteria, we prompt OpenAI’s o3 model to perform a comprehensive literature review of the relevant field.
-Each prompt includes a task description, example input-output pairs from the dataset, and instructions to generate a list of criteria considered important for performing the task – accompanied by reputable citations.
-We begin with this deep research approach to avoid over-reliance on any single expert’s perspective.
-Our goal is to synthesize insights from a broad array of books, journals, and academic publications to produce as comprehensive a list as possible.
-Step 2: Iteration with domain experts.
-To validate and improve the output from Step 1, we present the preliminary criteria list to a domain expert (see Figure 4 for details on each expert per domain).
-We ask the expert to (1) remove any incorrect or irrelevant criteria, (2) add any important ones that were missed, and (3) ensure that the list reflects a consensus that their peers would accept.
-The expert then refines the list until it accurately captures the field’s knowledge.
-An example criterion for sepsis classification is as follows: Advanced age (over 65 years) markedly increases susceptibility to rapid sepsis progression and higher mortality after infection.
-All Deep Research prompt templates and final expert alignment criteria lists for all domains are available in our GitHub repository.
+When answering knowledge-intensive questions like “Will this patient develop sepsis in the next 12 hours?” or “What kind of supernova produced these wavelengths?”, doctors and astrophysicists rely on domain-specific heuristics, prioritizing certain features over others based on their training and experience.
+For instance, in sepsis classification, an experienced clinician would typically emphasize features like advanced age and elevated temperature, while assigning lower importance to signals like patient demeanor, which are less directly indicative of sepsis risk.
+Thus, an LLM that makes the correct prediction using age and temperature is more expert-aligned than one that arrives at the same answer by focusing on only patient demeanor.
+We define the features that experts prioritize as expert alignment criteria.
+We draw inspiration from prior work (Yang et al., 2026; Li et al., 2025b), but seek to move beyond example-level rubrics to task-level criteria.
+Developing Expert Criteria.
+We work with five medical experts (surgeons and clinicians), three cosmologists (astrophysicists and astronomers), and two psychologists to develop the expert criteria for each of our datasets.
+See Figure 4 for details on collaborating experts.
+To save our domain experts’ time, we seed our initial list of expert criteria using o3 deep research (OpenAI, 2025c) to perform a comprehensive literature review of the relevant field.
+Each prompt includes a task description, examples, and instructions to generate a list of criteria considered important for performing the task – accompanied by reputable citations from academic books, journals, and conferences.
+Then, we present the preliminary criteria list to the relevant domain experts and ask them to (1) remove any incorrect or irrelevant criteria, (2) add any important ones that were missed, and (3) ensure that the list reflects a consensus among them.
+The experts then iteratively refine the list until they feel it accurately captures the field’s knowledge.
+An example criterion for sepsis classification is as follows: Advanced age (over 65 years) increases susceptibility to rapid sepsis progression and higher mortality after infection.
+All Deep Research prompt templates and final expert alignment criteria are available in our GitHub repository.
 3 T-FIX Pipeline
-Figure 4: Overview of datasets and domains in T-FIX.
-We evaluate LLM expert alignment across seven diverse domains, spanning cosmology, psychology, and medicine.
-For each dataset, we highlight the motivating task, input–output format, representative example, and the expert responsible for validating alignment criteria.
-The final row summarizes the expert alignment criteria used for scoring explanations in each domain.
-The column colors reflect dataset modality: blue indicates vision, yellow indicates language, and pink indicates time-series.
-LLM-generated explanations contain a mix of reasoning steps – some aligned with expert judgment, and others based on irrelevant information.
-To systematically evaluate such complex explanations, we first break them down into atomic claims, or standalone “features” that can be individually assessed for expert alignment.
-By scoring each feature separately and then aggregating these scores, we can compute an overall expert alignment score for the full explanation.
-See Figure 3 for an example of this multi-step process.
-Our T-FIX pipeline, built using GPT-4o, for evaluating expert alignment consists of three components:
+Next, we develop a framework to use these expert criteria to automatically evaluate explanations.
+LLM-generated explanations typically contain a mix of reasoning steps — some aligned with expert judgment, and others based on irrelevant information.
+To systematically evaluate such complex explanations, we first break them down into atomic claims, or standalone “features.” We then group these claims by expert criteria, score each feature group for expert alignment, and aggregate these scores to compute an overall alignment score for the full explanation.
+See Figure 3 for an example of this multi-step evaluation process.
+Our T-FIX pipeline for evaluating expert alignment consists of three components:
   1. 1.
@@
   2. 2.
-    Relevancy Filtering: Removing claims that are unsupported, speculative, or otherwise irrelevant to the model’s prediction.
+    Claim Filtration & Grouping: Removing claims that are irrelevant to the model’s prediction and grouping claims by relation to each expert criterion.
   3. 3.
-    Alignment Scoring: Measuring the degree of overlap between each remaining claim and domain expert criteria.
+    Alignment Scoring: Measuring how much each group of claims supports its corresponding expert criteria.
+We implement this pipeline using an evaluator LLM; results reported below use GPT-5-mini (OpenAI, 2025b).
 3.1 Stage 1: Atomic Claim Extraction
@@
 We achieve this by decomposing the explanation into atomic claims.
-An atomic claim is defined as a self-contained, indivisible statement that conveys a single verifiable fact, and can be fully understood without reference to the surrounding context.
-To extract atomic claims, we adapt prompting techniques from the claim decomposition literature wanner2024closer; gunjal2024molecularfactsdesideratadecontextualization and prompt GPT-4o to transform a free-form explanation into a list of fully decontextualized atomic claims.
+An atomic claim is defined as a self-contained, indivisible statement that conveys a single fact, and can be fully understood without reference to the surrounding context.
+To extract atomic claims, we adapt prompting techniques from the claim decomposition literature (Wanner et al., 2024; Gunjal and Durrett, 2024) and prompt GPT-5-mini to transform a free-form explanation into a list of fully decontextualized atomic claims.
 We treat each claim as representing a single “feature” in the LLM’s explanation.
-3.2 Stage 2: Relevancy Filtering
+3.2 Stage 2: Claim Filtering & Grouping
+Filtering.
 Not all extracted claims contribute meaningfully to expert reasoning.
-Some may be unsupported (i.e., references to content not present in the input), speculative (i.e., unfounded hypotheses), or otherwise irrelevant (e.g., repeating the model’s final prediction or citing unrelated information).
-Given domain experts’ preference for succinct, informative explanations, we prompt GPT-4o to remove such noisy claims by evaluating each atomic claim based on the original input.
-A claim is retained if it satisfies the following two criteria:
-(1) Clearly grounded in and supported by the input (i.e., not unfounded or speculative)
-(2) Directly contributes to explaining why the model made its prediction.
-On average, 72% of the claims generated in Stage 1 pass this relevancy filter and are carried forward for alignment scoring.
+Some may be unsupported (i.e., references to content not in the input), speculative (i.e., unfounded hypotheses), or otherwise irrelevant (e.g., repeating the model’s final prediction or citing unrelated information).
+Given domain experts’ preference for succinct, informative explanations, we prompt GPT-5-mini to remove such noisy claims by evaluating each atomic claim based on the original input.
+A claim is retained if it is (1) Clearly grounded in and supported by the input, and (2) Directly contributes to explaining why the model made its prediction.
+On average, 76.8% of the claims generated in Stage 1 pass this relevancy filter and are carried forward for grouping.
+Grouping.
+Next, we prompt GPT-5-mini to group extracted claims by their relationship to each expert criterion.
+For each criterion, the LLM selects the subset of claims that are topically related to that criterion.
+Formally, given a set of expert criteria $K=\{k_{1},\ldots,k_{m}\}$ and an explanation consisting of claims $C=\{c_{1},\ldots,c_{n}\}$, we define the claim group of criterion $k$ as:
+$G_{k}=\{c\in C:\mathsf{related}(c,k)\},$ (1)
+where we use GPT-5-mini to check the $\mathsf{related}$ predicate.
+Importance of Claim Grouping.
+We group claims at the criterion level because expert alignment is often determined by the joint content of multiple claims: claims can depend on one another, and some criteria are inherently compositional.
+For example, SIRS positivity in sepsis diagnosis requires meeting at least two symptoms from a predefined list, which may be expressed across several separate claims.
+Table 1: Interpretation of alignment labels in scoring atomic claims against expert criteria.
+Label Meaning
+complete The claims are all specific, directly relevant, and fully capture the meaning and intent of the expert criterion.
+partial The claims partially support the expert criterion but lack key details, use vague language, are overly general, or contain noise.
+none The claims reference something unrelated to the expert criterion, misinterprets the criterion’s meaning, or contradict what is stated in the criterion.
 3.3 Stage 3: Alignment Scoring
-In the final stage of our pipeline, we evaluate each retained atomic claim by comparing it to the domain-specific expert alignment criteria (see Section 2).
-This step quantifies how the reasoning in the LLM’s explanation reflects expert judgment.
-Given an atomic claim and a list of expert criteria, we prompt GPT-4o to measure the claim’s expert alignment in two steps:
-  1. 1.
-    Identify the most aligned expert criterion.
-The model selects the criterion whose focus and intent best match the core idea of the atomic claim.
-The model may also indicate that no criteria align with the claim.
-  2. 2.
-    Label claim alignment. The model scores how well the claim aligns with the chosen criterion: complete, partial, or none. See Table˜1 for details.
-Label Meaning
-none The claim references something unrelated to the expert category, or misinterprets the category’s meaning.
-partial The claim partially refers to the expert category but lacks key details, uses vague language, is overly general, or contains noise.
-complete The claim is specific, directly relevant, and fully captures the meaning and intent of the expert category.
-Table 1: Interpretation of alignment labels in scoring atomic claims against expert criteria.
+In the final stage of our pipeline, we score each criterion-level group of claims by assessing how well the group, taken as a whole, reflects its corresponding domain-specific expert criterion.
+That is, we wish to assign a numerical score to each claim group $G_{k}$.
+We compute this score by prompting the LLM GPT-5-mini to label the group’s alignment as complete, partial, or none (see Table˜1), and then assigning these labels to numerical values.
 For example, consider the expert criterion for sepsis classification: Advanced age (over 65 years).
 The claim ‘‘The patient is at risk as they are 72 years old’’ would be labeled complete, as it directly and fully supports the criterion.
-In contrast, the claim ‘‘The patient is at risk as they are 37’’ may be labeled partial: while it discusses patient age, the specific value does not align with the expert threshold for elevated risk.
-In contrast, the claim ‘‘The patient is NOT at risk as they are 37’’ would be labeled complete.
-Examples of claims with high and low alignment for each domain, along with rationale for why those scores were assigned, are provided in Table˜A4.
-3.4 Final Aggregation
-We assign an alignment label of none to the claims that were filtered out or did not align with any criteria.
-This ensures LLM-generated explanations are penalized for unsupported or speculative statements, irrelevant information, and misaligned reasoning.
-Then, we map alignment labels to numerical scores as follows: complete $\rightarrow 1$, partial $\rightarrow 0.5$, and none $\rightarrow 0$.
-We average these scores across all claims to produce a final expert alignment score for the explanation.
-The prompts for all three stages can be found in Appendix˜B and in our Github repository.
+The claim ‘‘The patient is NOT at risk as they are 37’’ would also be labeled complete.
+In contrast, the claim ‘‘The patient is at risk as they are 64’’ may be labeled partial: It reasons about patient age correctly, but the specific value does not align with the expert threshold for elevated risk.
+We express the LLM-based labeling of a claim group with respect to its expert criterion, $\mathcal{L}(G_{k},k)\in\{\textit{complete},\textit{partial},\textit{none}\}$, as the group-wise alignment:
+[…]
+Examples of claims with high and low alignment for each domain, along with rationale for why those scores were assigned, are provided in Table˜A5.
+3.4 Final Aggregation: T-FIXscore
+Now that we’ve measured the alignment of each claim group, we turn to the harder task of isolating the alignment score of each individual claim.
+We calculate claim-level alignment scores by taking the maximum score across all groups that contain the claim.
+Any claim that supports at least one criterion is thus treated as aligned with expert intuition.
+For the set of claims $C$, criteria $K$, and the corresponding criterion-level claim groups $\{G_{1},\dots,G_{m}\}$, we formally define claim alignment as:
+$\mathsf{CAlign}(c_{i},K)=\max_{k\in K:c_{i}\in G_{k}}\big(\mathsf{GAlign}(G_{k},k)\big)$ (3)
+where we let the maximum over the empty set be 0, such as in the case where a claim is unrelated to any criterion. This penalizes irrelevant information and misaligned reasoning.
+Finally, we average the per-claim alignment score to compute the final explanation alignment score as:
+$\mathsf{T\textrm{-}FIXscore}(C,K)=\frac{1}{n}\sum_{i=1}^{n}\mathsf{CAlign}(c_{i},K)$ (4)
+We draw inspiration from Jin et al.
+(2025), which details a method for aggregating expert alignment for input feature groups typically extracted from trained classification models.
+With $\mathsf{T\textrm{-}FIXscore}$, we extend the metric that Jin et al.
+(2025) proposes to free-form text explanations from generative models not directly linked to explicit feature groups.
+3.5 Prompt Development with Expert Feedback
+We met with our domain experts throughout the prompt development process for each step of our pipeline.
+We asked what they would do differently from the LLM, incorporated their feedback, and iterated on all prompts multiple times, until the experts were satisfied with the results.
+Our prompt iteration process is detailed in Appendix F. The final prompts for all three stages can be found in Appendix˜E and in our Github repository.
 4 Pipeline Validation
-Given our pipeline relies on multiple curated GPT-4o prompts, we want to ensure that the extracted and filtered claims are accurate, and that the final alignment scores match domain expert intuition.
-To do this, we conduct a quantitative evaluation via annotation, and a qualitative evaluation via domain expert review.
+Given our pipeline relies on multiple curated GPT-5-mini prompts, we want to ensure that the extracted and filtered claims are accurate, and that the final alignment scores match domain expert intuition.
+To do this, we conduct a quantitative evaluation via an annotation study, and a qualitative evaluation via domain expert interviews.
 Annotation study.
@@
 Six annotators participate, with two annotators per example.
-[^2] Annotators are PhD students in machine learning at an American university with experience evaluating LLM outputs.
-Annotators evaluate each stage using a unified labeling scheme: (A) fully accurate, (B) partially accurate, or (C) incorrect.
-For atomic claim extraction, they review the explanation and its extracted claims, labeling outputs as (A) correct decomposition, (B) 1–2 claims missing/incorrect, or (C) 3+ claims missing/incorrect.
-For relevancy filtering, they assess whether claims were appropriately kept or removed, labeling results as (A) all correct, (B) 1–2 incorrect, or (C) 3+ incorrect.
-For expert alignment scoring, annotators review the filtered claims and expert categories, labeling each as (A) matches expert judgment, (B) borderline, or (C) does not match expert judgment.
-These labels are then mapped to accuracy scores of 1.0, 0.5, and 0.0, respectively.
-Table˜2 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX domains, indicating moderate-to-substantial agreement and supporting the reliability of our pipeline.
+[^1] Annotators are PhD students in machine learning trained extensively by relevant domain experts to evaluate LLM outputs.
+See Appendix B for details on study setup.
+Table˜A1 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX datasets, indicating substantial agreement and supporting the reliability of our pipeline.
 Domain-level metrics appear in Table˜A3.
-Domain expert interview.
-To qualitatively validate the expert alignment scores in Stage 3 of our pipeline, we interview each domain expert and have them analyze 10-20 claims for their dataset, discussing noticed successes and failures.
+Domain expert interviews.
+To qualitatively validate the expert alignment scores in Stage 3 of our pipeline, we conduct ongoing interviews with each domain expert and have them annotate and analyze 10-20 claims for their dataset, discussing noticed successes and failures.
 Across domains, experts largely agreed with the LLM alignment labels, confirming that the evaluation framework reliably captures expert reasoning.
-Agreement was highest in domains with well-defined interpretive criteria (e.g., physics and psychology) and somewhat lower in medicine, where decomposing multi-symptom medical statements occasionally caused partial rather than full alignment.
+Agreement was highest in domains with well-defined interpretive criteria (e.g., physics and psychology).
+Additional prompt iteration incorporating interview feedback was required to achieve high agreement in medicine, as multi-symptom, compositional medical statements were more complex.
 Overall, expert validation supports the interpretability and trustworthiness of the T-FIX evaluation pipeline.
-We provide a summary of each domain expert’s interview in Appendix˜E.
-Pipeline Stage $\mathcal{N}$ Accuracy Cohen’s $\kappa$
-Claim Extraction 35 0.943 0.717
-Relevancy Filtering 295 0.871 0.402
-Expert Alignment 211 0.923 0.405
-Table 2: Pipeline validation: Average accuracy for $\mathcal{N}$ samples across all T-FIX domains and annotator agreement – Cohen’s $\kappa$ for each stage in our pipeline.
-Domain-specific statistics are provided in Table˜A3.
+We provide a summary of each domain expert’s interview in Appendix˜D.
+Figure 4: Overview of datasets and domains in T-FIX.
+We evaluate LLM expert alignment across seven diverse scientific tasks, spanning cosmology, psychology, and medicine.
+For each dataset, we highlight the motivation, I/O format, representative example, and the key expert responsible for validation.
+The final row summarizes the expert alignment criteria used for scoring explanations in each domain.
+The column colors reflect dataset modality: blue indicates vision, yellow indicates language, and pink indicates time-series.
 5 Included Datasets
+Table 2: Evaluating top LLMs on T-FIX using GPT-5-mini.
+We report the average $\mathsf{T\textrm{-}FIXscore}$ across all examples in the dataset.
+Corresponding accuracies are in Table˜A2 and baseline prompting strategies are described in Section 6.
+We report $\pm$ standard deviation over 1000 bootstrap samples.
+In general, more sophisticated prompting strategies do not improve expert alignment of the explanations.
 Cosmology Psychology Medicine
 Baseline Mass Maps Supernova Politeness Emotion Cholecystectomy Cardiac Sepsis
-GPT-4o (openai2024gpt4ocard)
-Vanilla 0.334 0.726 0.564 0.494 0.236 0.427 0.535
-Chain-of-Thought 0.320 0.697 0.534 0.514 0.278 0.472 0.541
-Socratic Prompting 0.329 0.708 0.536 0.501 0.299 0.466 0.547
-SubQ Decomposition 0.279 0.719 0.523 0.439 0.294 0.435 0.553
-o1 (openai2024openaio1card)
-Vanilla 0.528 0.570 0.560 0.521 0.370 0.446 0.525
-Chain-of-Thought 0.501 0.552 0.562 0.555 0.387 0.428 0.560
-Socratic Prompting 0.428 0.575 0.502 0.397 0.371 0.370 0.555
-SubQ Decomposition 0.430 0.582 0.482 0.480 0.328 0.398 0.574
-Gemini-2.0-Flash (google2024gemini2flash)
-Vanilla 0.454 0.650 0.552 0.495 0.335 0.445 0.544
-Chain-of-Thought 0.442 0.666 0.505 0.489 0.312 0.465 0.539
-Socratic Prompting 0.223 0.666 0.490 0.470 0.323 0.409 0.560
-SubQ Decomposition 0.323 0.660 0.509 0.449 0.325 0.408 0.557
-Claude-3.5-Sonnet (anthropic2025claude35)
-Vanilla 0.622 0.581 0.558 0.533 0.211 0.459 0.564
-Chain-of-Thought 0.605 0.604 0.565 0.493 0.228 0.440 0.563
-Socratic Prompting 0.616 0.613 0.529 0.474 0.235 0.448 0.557
-SubQ Decomposition 0.532 0.594 0.554 0.503 0.221 0.460 0.530
-Table 3: Evaluating top LLMs on T-FIX.
-We report the average expert alignment score across all examples in the dataset.
-Corresponding accuracies are in Table˜A2 and baseline prompting strategies are described in Section 6.
-In general, more sophisticated prompting strategies do not necessarily improve expert alignment of explanation.
-T-FIX contains seven open-source datasets, spanning the fields of cosmology, psychology, and medicine.
+GPT-5.2-Pro (OpenAI, 2025a)
+Vanilla $0.901_{\pm 0.01}$ $0.845_{\pm 0.01}$ $0.882_{\pm 0.01}$ $0.906_{\pm 0.01}$ $0.538_{\pm 0.03}$ $0.539_{\pm 0.02}$ $0.425_{\pm 0.03}$
+[…]
+Socratic Prompting $0.804_{\pm 0.01}$ $0.714_{\pm 0.01}$ $0.763_{\pm 0.02}$ $0.726_{\pm 0.02}$ $0.656_{\pm 0.02}$ $0.488_{\pm 0.02}$ $0.381_{\pm 0.02}$
+SubQ Decomposition $0.812_{\pm 0.01}$ $0.727_{\pm 0.01}$ $0.857_{\pm 0.01}$ $0.774_{\pm 0.02}$ $0.639_{\pm 0.02}$ $0.489_{\pm 0.02}$ $0.356_{\pm 0.02}$
+GPT-5-mini (OpenAI, 2025b)
+[…]
+Chain-of-Thought $0.759_{\pm 0.01}$ $0.876_{\pm 0.01}$ $0.767_{\pm 0.02}$ $0.830_{\pm 0.02}$ $0.609_{\pm 0.03}$ $0.507_{\pm 0.02}$ $0.395_{\pm 0.03}$
+Socratic Prompting $0.729_{\pm 0.01}$ $0.781_{\pm 0.01}$ $0.641_{\pm 0.02}$ $0.690_{\pm 0.02}$ $0.506_{\pm 0.03}$ $0.444_{\pm 0.02}$ $0.399_{\pm 0.02}$
+SubQ Decomposition $0.720_{\pm 0.02}$ $0.838_{\pm 0.01}$ $0.809_{\pm 0.02}$ $0.846_{\pm 0.02}$ $0.528_{\pm 0.03}$ $0.443_{\pm 0.01}$ $0.419_{\pm 0.03}$
+Claude-Opus-4.5 (Anthropic, 2025b)
+[…]
+[…]
+Socratic Prompting $0.835_{\pm 0.01}$ $0.683_{\pm 0.01}$ $0.802_{\pm 0.02}$ $0.789_{\pm 0.02}$ $0.647_{\pm 0.01}$ $0.653_{\pm 0.02}$ $0.390_{\pm 0.02}$
+SubQ Decomposition $0.765_{\pm 0.01}$ $0.736_{\pm 0.01}$ $0.800_{\pm 0.02}$ $0.821_{\pm 0.02}$ $0.546_{\pm 0.01}$ $0.620_{\pm 0.02}$ $0.386_{\pm 0.02}$
+Claude-Haiku-4.5 (Anthropic, 2025a)
+Vanilla $0.742_{\pm 0.01}$ $0.630_{\pm 0.02}$ $0.873_{\pm 0.02}$ $0.881_{\pm 0.02}$ $0.677_{\pm 0.01}$ $0.529_{\pm 0.02}$ $0.355_{\pm 0.02}$
+Chain-of-Thought $0.781_{\pm 0.01}$ $0.640_{\pm 0.02}$ $0.871_{\pm 0.02}$ $0.887_{\pm 0.02}$ $0.722_{\pm 0.01}$ $0.510_{\pm 0.02}$ $0.368_{\pm 0.02}$
+Socratic Prompting $0.686_{\pm 0.01}$ $0.658_{\pm 0.02}$ $0.838_{\pm 0.01}$ $0.842_{\pm 0.01}$ $0.687_{\pm 0.01}$ $0.492_{\pm 0.02}$ $0.300_{\pm 0.02}$
+SubQ Decomposition $0.626_{\pm 0.02}$ $0.537_{\pm 0.02}$ $0.769_{\pm 0.02}$ $0.841_{\pm 0.02}$ $0.616_{\pm 0.01}$ $0.505_{\pm 0.02}$ $0.308_{\pm 0.02}$
+Gemini-2.5-Pro (team, 2025)
+Vanilla $0.801_{\pm 0.01}$ $0.715_{\pm 0.02}$ $0.843_{\pm 0.02}$ $0.888_{\pm 0.02}$ $0.697_{\pm 0.01}$ $0.315_{\pm 0.02}$ $0.511_{\pm 0.02}$
+Chain-of-Thought $0.812_{\pm 0.01}$ $0.697_{\pm 0.02}$ $0.877_{\pm 0.02}$ $0.873_{\pm 0.02}$ $0.691_{\pm 0.02}$ $0.264_{\pm 0.02}$ $0.525_{\pm 0.02}$
+Socratic Prompting $0.671_{\pm 0.01}$ $0.616_{\pm 0.02}$ $0.682_{\pm 0.02}$ $0.655_{\pm 0.02}$ $0.670_{\pm 0.01}$ $0.281_{\pm 0.02}$ $0.466_{\pm 0.02}$
+SubQ Decomposition $0.731_{\pm 0.01}$ $0.635_{\pm 0.02}$ $0.759_{\pm 0.02}$ $0.836_{\pm 0.02}$ $0.654_{\pm 0.01}$ $0.221_{\pm 0.01}$ $0.522_{\pm 0.02}$
+Gemini-2.5-Flash (team, 2025)
+[…]
+[…]
+Socratic Prompting $0.790_{\pm 0.01}$ $0.683_{\pm 0.02}$ $0.831_{\pm 0.02}$ $0.812_{\pm 0.02}$ $0.620_{\pm 0.01}$ $0.315_{\pm 0.02}$ $0.484_{\pm 0.02}$
+SubQ Decomposition $0.800_{\pm 0.01}$ $0.688_{\pm 0.02}$ $0.816_{\pm 0.02}$ $0.841_{\pm 0.02}$ $0.688_{\pm 0.01}$ $0.300_{\pm 0.02}$ $0.506_{\pm 0.02}$
+T-FIX aggregates seven open-source datasets, spanning the fields of cosmology, psychology, and medicine.
 To assess LLM explanations across multiple modalities, we include text, vision, and time-series datasets.
 We select these seven datasets due to the availability of domain experts willing to work with us for these tasks.
-As running T-FIX requires querying LLMs, many of which follow a pay-as-you-go API structure, we cap the benchmark at 700 (100 per dataset) in order for T-FIX to be accessible to as many researchers as possible.
-We select a subset of 100 examples from the test set of each open-source dataset in T-FIX, and balance this sampling across classes when possible.
+As running T-FIX requires querying LLMs, many of which follow a pay-as-you-go API structure, we cap the benchmark at 700 (100 per dataset) in order to manage costs and enable T-FIX to be accessible to as many researchers as possible.
+We select 100 examples from the test set of each open-source dataset in T-FIX, and balance this sampling across classes when possible.
 We provide an overview of the included open-source datasets in Figure 4.
-See Appendix˜D for additional details about the motivation, task, and prompting procedure for each dataset.
+See Appendix˜G for additional details about the motivation, task, and prompting procedure for each dataset.
 6 Experiments
@@
   2. 2.
-    Chain-of-Thought (CoT): The LLM is prompted to reason step-by-step through intermediate steps before answering, supporting more accurate responses on multi-step tasks.
+    Chain-of-Thought (CoT): The LLM is prompted to reason step-by-step through intermediate steps before answering, improving responses on multi-step tasks.
   3. 3.
@@
   4. 4.
-    Subquestion Decomposition: The LLM is guided to break down a complex task into simpler subquestions, answer them individually, and then synthesize a final response.
-Domain-specific prompts are in Appendix˜D, with templates for the above prompting strategies in Figure˜A5.
-Results for GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet  [^3] We only select LLMs with vision support and context windows long enough to accommodate our time-series datasets.
-All models are accessed in September 2025.
-are in Table˜3.
-Role of LLM in pipeline.
+    Subquestion Decomposition: The LLM is guided to break down the task into simpler subquestions, answer them individually, and then synthesize a final response.
+Domain-specific prompts are in Appendix˜G, with templates for the above prompting strategies in Figure˜A6.
+Results for GPT-5.2, GPT-5-mini, Gemini-2.5-Flash, Gemini-2.5-Pro, Claude-Opus-4.5, and Claude-Haiku-4.5222We only select stable LLMs (i.e., not in preview mode) with vision support and context windows long enough to accommodate our time-series datasets.
+All models are accessed in January 2026.
+are in Table˜2.
+Role of evaluator LLM.
 Because our prompts are tightly constrained and refined through multiple iterations, the evaluation pipeline is largely robust to the choice of LLM we use.
-We validate this by comparing results across three evaluator models — GPT-4o, Gemini-2.0-Flash, and Qwen2.5-VL-7B-Instruct — on one dataset per domain.
-Despite minor variations, rankings and alignment score ranges remain consistent, confirming that the pipeline is robust to evaluator choice.
-Further details are provided in Appendix˜C.
+We validate this by comparing results across three evaluator models: GPT-5-mini, Gemini-2.5-Flash-Lite, and Qwen2.5-VL-7B-Instruct, on one dataset per domain.
+Despite minor variations, rankings and alignment score ranges remain consistent, confirming that the pipeline is robust to evaluator choice. Further details are provided in Appendix˜C.
 7 Analysis
-In this section, we analyze how LLMs distribute reasoning across expert criteria and whether higher task accuracy indicates better expert alignment.
-Figure 5: Shannon Entropy of expert alignment criteria for GPT-4o.
+In this section, we analyze how LLMs distribute reasoning across expert criteria and whether higher expert alignment indicates higher accuracy.
+Figure 5: Shannon Entropy of expert alignment criteria for GPT-5-mini.
 For each prompting baseline, we show coverage of each domain’s explanations across all expert criteria – a high value indicates the LLM considers many criteria across examples, while a low value indicates the LLM focuses on the same criteria repeatedly.
-Figure 6: Expert Alignment vs.
-Accuracy Correlation Heatmap, averaged across GPT-4o, o1, Gemini-2.0-Flash, and Claude-3.5-Sonnet.
-Blue indicates positive correlation, red is negative, gray is no correlation.
+Figure 6: Expert Alignment vs. Accuracy Correlation Heatmap, averaged across all 6 models in Table˜2. Blue indicates positive correlation, red is negative, gray is no correlation. $^{*}p<0.05$.
 7.1 Coverage of Expert Alignment Criteria
@@
 Instead, we analyze coverage at the dataset level – whether different prompting strategies lead to a broader utilization of expert criteria across all questions within a domain.
-Figure 5 presents the Shannon entropy of GPT-4o’s covered expert criteria in each domain.
-We observe a correlation between entropy and performance: domains where GPT-4o underperforms (e.g., Cholecystectomy, Supernova) show lower entropy, indicating limited criteria coverage.
-In contrast, well-performing domains (e.g., Politeness, Sepsis) exhibit more uniform coverage, equally taking into account all expert criteria.
-This suggests that LLMs that reason uniformly over expert alignment criteria perform better – a promising insight for future work in prompting or training models to incorporate a broader range of expert reasoning.
+Figure 5 presents the Shannon entropy of GPT-5-mini’s covered expert criteria in each domain.
+We observe a correlation between entropy and performance: domains where GPT-5-mini underperforms (e.g., Cholecystectomy, Supernova) show lower entropy, indicating limited criteria coverage.
+In contrast, higher-performing domains (e.g., Politeness, Sepsis) exhibit more uniform coverage, equally taking into account all expert criteria.
+This suggests that LLMs that reason uniformly over expert alignment criteria perform better — a promising insight for future work in prompting or training models to incorporate a broader range of expert reasoning.
 7.2 Expert Alignment vs. Accuracy
-T-FIX focuses on evaluating explanation quality, but we are also interested in understanding the relationship between expert alignment and prediction accuracy.
-Specifically, we ask: Does higher answer accuracy correspond to stronger expert alignment?
-Figure 6 shows the Pearson correlation of expert alignment (see Table˜3) with accuracy (see Table˜A2) for each domain, averaged across models.
-In some domains with higher performance, like Cholecystectomy and Emotion, we do observe higher expert alignment as well.
-However, the overall correlation is weak across domains.
+T-FIX evaluates explanation quality, but we also analyze how expert alignment relates to prediction accuracy. Specifically, we ask: Does higher answer accuracy correspond to stronger expert alignment?
+Figure 6 shows Pearson correlations between expert alignment (Table˜2) and accuracy (Table˜A2), averaged across models.
+Across domains, the vast majority of correlations are statistically insignificant or low in magnitude.
 The heatmap suggests today’s high-performing LLMs do not appear to rely on expert reasoning.
-Future research is needed to explore whether aligning model reasoning with expert criteria – via training objectives or prompting – can improve downstream performance.
+Future research is needed to explore whether explicitly aligning model reasoning with expert criteria — via training objectives or prompting — can improve downstream performance.
 8 Related Work
-Evaluating LLM explanations.
-Common explanation methods for LLMs include feature attribution (e.g., LIME, SHAP ribeiro2016lime; lundberg2017shap), counterfactuals, and self-generated explanations im2023evaluatingutilitymodelexplanations; zhao2023explainabilitylargelanguagemodels.
-Some models are also trained to produce human-readable justifications camburu2018esnli.
-To assess explanation quality and utility, recent work highlights criteria such as faithfulness (alignment with the model’s reasoning) and plausibility (how convincing it is to humans) jacovi2020faithfulness; zhou2021evaluating; agarwal2024faithfulness.
-Human studies show mixed outcomes: explanations sometimes aid understanding hase2020evaluating; bansal2021whole, but can also offer little value or cause over-trust wang2023fair.
-A promising alternative is to use LLMs as automatic judges of explanation quality zheng2023judging; chen2024judge, providing a scalable substitute for expensive human evaluation; we adopt this approach in T-FIX.
+Evaluating LLM explanations at scale.
+Common methods include self-generated explanations (Im et al., 2023; Zhao et al., 2023) or human-readable justifications (Camburu et al., 2018).
+To assess explanation quality and utility, faithfulness is analyzed (Jacovi and Goldberg, 2020; Zhou et al., 2021; Agarwal et al., 2024).
+Human studies show mixed outcomes: explanations sometimes aid understanding (Hase and Bansal, 2020; Bansal et al., 2021), but can also offer little value or cause over-trust (Wang et al., 2023).
+A promising alternative is to use LLMs as automatic judges of explanation quality (Zheng et al., 2023; Chen et al., 2024), providing a scalable substitute for expensive human evaluation; we adopt this approach in T-FIX.
 Domain & expert alignment.
-Concept-based models constrain parts of the network to predict high-level, human-defined concepts, enabling incorporation of domain knowledge into final predictions koh2020concept.
-Extensions of concept bottlenecks and related methods aim to align latent representations with semantically meaningful features kim2018tcav; chen2020concept; ghorbani2019ace, potentially grouped for expert interpretability jin2024fix.
-In NLP, integrating human knowledge has included collecting human-written explanation datasets to train models camburu2018esnli and using learned explanations to guide predictions bhatt2020explainable.
-To our knowledge, no prior work explicitly evaluates text explanations for expert alignment like T-FIX.
+Prior work incorporates domain knowledge through concept-based representations that constrain or structure predictions (Koh et al., 2020; Kim et al., 2018; Chen et al., 2020; Ghorbani et al., 2019; Jin et al., 2025), and through explanation supervision in NLP (Camburu et al., 2018; Bhatt et al., 2020).
+Existing efforts to evaluate expert alignment are intensive and small-scale (Kutbi et al., 2025; Hyk et al., 2025)), while larger domain benchmarks rarely assess expert alignment beyond validation (e.g., Astro-QA (Li et al., 2025a), CounselBench (Li et al., 2025b), Mentalchat16k (Xu et al., 2025)) and rubric-based medical evaluations remain tied to fixed examples (Yang et al., 2026).
+To our knowledge, no prior work explicitly evaluates expert alignment like T-FIX does.
 9 Conclusion
 We introduce T-FIX, the first benchmark designed to evaluate LLM explanations for expert alignment across seven knowledge-intensive domains.
-Our analysis reveals that today’s models struggle to generate explanations that experts would rely on, highlighting a critical area for improvement.
+Unlike existing evaluation frameworks, T-FIX front-loads expert effort into a reusable framework that enables automatic expert-alignment scoring of unseen explanations.
 Future work may include exploring instruction-tuning LLMs to generate explanations with strong expert alignment, extending T-FIX to additional domains, and Human-Computer Interaction studies exploring how expert-aligned explanations affect real-world decision-making by practitioners.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Limitations
@@
 We hope the issue of prompt dependency diminishes with future models that are more robust and less susceptible to tiny prompt ablations.
-While our evaluation pipeline currently uses GPT-4o for scoring, it is model-agnostic by design, and we encourage future work to apply or adapt the pipeline with other LLMs to improve robustness and reduce evaluator-model entanglement.
+While our evaluation pipeline currently uses GPT-5-mini for scoring, it is model-agnostic by design, and we encourage future work to apply or adapt the pipeline with other LLMs to improve robustness and reduce evaluator-model entanglement.
 For pipeline validation, we conduct a user study where we annotate 35 examples.
@@
 Future work should also involve domain experts vetting the pipeline in addition to recruited annotators.
-In addition, we only have one expert to validate the expert alignment criteria for each domain.
-Although our usage of a deep research LLM minimizes over-reliance on a single domain expert, multiple experts would have been better to create the expert criteria.
+In addition, we only have a small number of experts to validate the expert alignment criteria for each domain.
+Having additional experts would have been better to create a more robust expert criteria.
 We were constrained by the availability of domain experts.
-Our experiments focus on a set of four models and four prompting strategies, and including additional models and strategies could provide a more comprehensive set of baseline results.
-Though many other high-performing LLMs and prompting techniques exist as of September 2025, we are conscious of budget and the environmental impact of running multiple experiments using T-FIX.
-Ethical Considerations
-Using LLMs in the domains we describe in T-FIX, especially those relating to medicine, poses a unique set of risks and challenges.
-We do not advocate that LLMs should replace domain experts in these tasks; rather, T-FIX should serve as a step towards experts being able to use LLMs in a reliable and trustworthy way.
+Our experiments focus on a set of six models and four prompting strategies, and including additional models and strategies could provide a more comprehensive set of baseline results.
+Though many other high-performing LLMs and prompting techniques exist as of January 2026, we are conscious of budget and the environmental impact of running multiple experiments using T-FIX.
 Additionally, LLMs are constantly changing, especially those that are company-owned and not open-source.
 This poses potential issues relating to the reproducibility of our baseline results as time progresses and advances are made.
-Lastly, nearly all LLMs contain biases – some harmful – that may propagate up in a system built off of these models. All users of T-FIX must be conscious of this risk.
-Evaluation Model
-GPT-4o
-Gemini-2.0-Flash
-Qwen2.5-VL-7B-Instruct
-(Qwen2.5-VL)
-Mass Maps
-GPT-4o 0.334 0.394 0.507
-o1 0.528 0.581 0.596
-Gemini-2.0-Flash 0.454 0.483 0.431
-Claude-3.5-Sonnet 0.622 0.692 0.651
-Emotion
-GPT-4o 0.494 0.595 0.579
-o1 0.521 0.632 0.567
-Gemini-2.0-Flash 0.495 0.621 0.567
-Claude-3.5-Sonnet 0.533 0.610 0.565
-Cholecystectomy
-GPT-4o 0.236 0.000 0.145
-o1 0.370 0.000 0.234
-Gemini-2.0-Flash 0.335 0.000 0.153
-Claude-3.5-Sonnet 0.211 0.000 0.146
-Table A1: Model ablation study for evaluation pipeline.
-Across domains, model rankings are largely consistent across evaluators, indicating the robustness of our evaluation pipeline.
-Comparison of base models (explanation generator LLMs) across evaluator models (pipeline LLMs) for one dataset per domain.
-Best per evaluator is in bold; second-best is underlined.
-Cosmology Psychology Medicine
-Baseline Mass Maps Supernova Politeness Emotion Cholecystectomy Cardiac Sepsis
-GPT-4o
-Vanilla 0.039 0.103 0.916 0.259 0.075 0.567 0.657
-Chain-of-Thought 0.044 0.093 0.824 0.286 0.103 0.460 0.714
-Socratic Prompting 0.044 0.127 0.829 0.277 0.115 0.462 0.657
-SubQ Decomposition 0.049 0.118 0.837 0.304 0.115 0.485 0.657
-o1
-Vanilla 0.044 0.170 0.784 0.304 0.194 0.656 0.752
-Chain-of-Thought 0.045 0.146 0.818 0.339 0.177 0.685 0.750
-Socratic Prompting 0.042 0.155 0.793 0.348 0.155 0.646 0.755
-SubQ Decomposition 0.044 0.147 0.818 0.321 0.138 0.695 0.780
-Gemini-2.0-Flash
-Vanilla 0.045 0.145 0.831 0.223 0.253 0.577 0.654
-Chain-of-Thought 0.042 0.118 0.837 0.232 0.255 0.558 0.663
-Socratic Prompting 0.041 0.118 0.809 0.232 0.159 0.592 0.661
-SubQ Decomposition 0.053 0.109 0.773 0.241 0.249 0.562 0.688
-Claude-3.5-Sonnet
-Vanilla 0.053 0.127 0.962 0.241 0.146 0.485 0.709
-Chain-of-Thought 0.050 0.118 1.012 0.268 0.150 0.538 0.735
-Socratic Prompting 0.044 0.118 0.998 0.232 0.145 0.508 0.748
-SubQ Decomposition 0.050 0.136 0.990 0.259 0.149 0.485 0.741
+Finally, using LLMs in the domains we describe in T-FIX, especially those relating to medicine, poses a unique set of risks and challenges.
+We do not advocate that LLMs should replace domain experts in these tasks; rather, T-FIX should serve as a step towards experts being able to use LLMs in a reliable and trustworthy way.
+Appendix A Extending T-FIX to a New Domain
+Though T-FIX covers a wide range of knowledge-intensive settings, it can easily be extended to additional domains.
+A key contribution of the T-FIX benchmark is the framework: we create a pipeline to score any free-form text explanation for expert alignment given a set of expert criteria.
+Additionally, we iterate extensively on all our prompt templates to ensure all T-FIX users need to do is input their task-specific details and perform no additional prompt engineering for good results.
+To add a new domain to T-FIX, we advise you to follow these steps:
+  1. 1.
+    Generate criteria: Use the deep research prompt template shown in Figure˜A5 to generate a list of expert alignment criteria for your domain.
+Optionally, have a domain expert vet the generated criteria.
+  2. 2.
+    Modify prompts: Modify the prompt templates outlined in Figure˜A1, Figure˜A2, Figure˜A3, and Figure˜A4 with your task description, few-shot examples, and generated expert criteria.
+  3. 3.
+    Run T-FIX: Plug in your prompts for each stage of the pipeline and run T-FIX on your dataset!
+We encourage you to contact the authors of this work if you need additional assistance setting up your custom domain.
+Appendix B Annotation Study Setup
+Annotator qualifications.
+Annotators are PhD students in machine learning and the authors of this paper.
+Each author annotating for a scientific domain has had multiple meetings with that domain’s experts, and been taught exactly how to evaluate an LLM’s explanation for expert alignment.
+Note that this annotation study is meant to support the output of the T-FIX pipeline in conjunction with the qualitative study described in Appendix D.
+Annotators evaluate each stage using a unified labeling scheme: (A) fully accurate, (B) partially accurate, or (C) incorrect.
+For atomic claim extraction, they review the explanation and its extracted claims, labeling outputs as (A) correct decomposition, (B) 1–2 claims missing/incorrect, or (C) 3+ claims missing/incorrect.
+For filtering, they assess whether claims were appropriately kept or removed, labeling results as (A) all correct, (B) 1–2 incorrect, or (C) 3+ incorrect.
+For expert alignment scoring, annotators review the grouped claims and expert criteria, labeling each as (A) consistent with expert-defined criteria, (B) borderline, or (C) inconsistent with expert-defined criteria.
+These labels are then mapped to accuracy scores of 1.0, 0.5, and 0.0, respectively.
+Table˜A1 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX datasets, indicating moderate-to-substantial agreement and supporting the reliability of our pipeline.
+Domain-level metrics appear in Table˜A3.
+Table A1: Pipeline validation: Average accuracy for $\mathcal{N}$ samples across all T-FIX domains and annotator agreement – Cohen’s $\kappa$ for each stage in our pipeline.
+Domain-specific statistics are provided in Table˜A3.
+Pipeline Stage $\mathcal{N}$ Accuracy Cohen’s $\kappa$
+Claim Extraction 35 0.943 0.717
+Filtering & Grouping 295 0.871 0.402
+Expert Alignment 211 0.923 0.405
 Table A2: Evaluating top LLMs on T-FIX.
@@
 Baseline implementations are described in Section 6.
+We report $\pm$ standard deviation over 1000 bootstrap samples.
+Cosmology Psychology Medicine
+Baseline Mass Maps Supernova Politeness Emotion Cholecystectomy Cardiac Sepsis
+GPT-5.2-Pro (OpenAI, 2025a)
+Vanilla $0.060^{*}_{\pm 0.01}$ $0.236_{\pm 0.04}$ $0.801^{*}_{\pm 0.11}$ $320_{\pm 0.05} 0.0$$014^{*}_{\pm 0.01}$ $0.612_{\pm 0.04}$ $0.627_{\pm 0.05}$
+Chain-of-Thought $0.058^{*}_{\pm 0.01}$ $0.218_{\pm 0.04}$ $0.868^{*}_{\pm 0.12}$ $0.340_{\pm 0.05}$ $0.162^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.618_{\pm 0.05}$
+Socratic Prompting $0.062^{*}_{\pm 0.01}$ $0.218_{\pm 0.04}$ $0.841^{*}_{\pm 0.12}$ $300_{\pm 0.05} 0.0$$098^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.637_{\pm 0.05}$
+SubQ Decomposition $0.061^{*}_{\pm 0.01}$ $0.227_{\pm 0.04}$ $0.871^{*}_{\pm 0.12}$ $0.340_{\pm 0.05}$ $0.146^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.627_{\pm 0.05}$
+GPT-5-mini (OpenAI, 2025b)
+Vanilla $0.048^{*}_{\pm 0.00}$ $0.191_{\pm 0.04}$ $1.004^{*}_{\pm 0.15}$ $0.410_{\pm 0.05}$ $0.149^{*}_{\pm 0.01}$ $0.585_{\pm 0.04}$ $0.624_{\pm 0.05}$
+Chain-of-Thought $0.046^{*}_{\pm 0.00}$ $0.164_{\pm 0.04}$ $1.048^{*}_{\pm 0.14}$ $350_{\pm 0.05} 0.0$$067^{*}_{\pm 0.01}$ $0.577_{\pm 0.04}$ $0.676_{\pm 0.05}$
+Socratic Prompting $0.047^{*}_{\pm 0.00}$ $0.200_{\pm 0.04}$ $1.018^{*}_{\pm 0.15}$ $310_{\pm 0.05} 0.0$$018^{*}_{\pm 0.01}$ $0.585_{\pm 0.04}$ $0.627_{\pm 0.05}$
+SubQ Decomposition $0.050^{*}_{\pm 0.00}$ $0.164_{\pm 0.03}$ $1.024^{*}_{\pm 0.15}$ $340_{\pm 0.05} 0.0$$011^{*}_{\pm 0.00}$ $0.577_{\pm 0.04}$ $0.608_{\pm 0.05}$
+Claude-Opus-4.5 (Anthropic, 2025b)
+Vanilla $035^{*}_{\pm 0.00} 0.0$$082_{\pm 0.03}$ $0.701^{*}_{\pm 0.08}$ $0.300_{\pm 0.05}$ $0.273^{*}_{\pm 0.01}$ $0.454_{\pm 0.05}$ $0.618_{\pm 0.05}$
+Chain-of-Thought $0.036^{*}_{\pm 0.00}$ $0.118_{\pm 0.03}$ $0.814^{*}_{\pm 0.12}$ $0.300_{\pm 0.05}$ $0.287^{*}_{\pm 0.01}$ $0.469_{\pm 0.04}$ $0.634_{\pm 0.05}$
+Socratic Prompting $0.035^{*}_{\pm 0.00}$ $0.109_{\pm 0.03}$ $0.708^{*}_{\pm 0.08}$ $0.310_{\pm 0.05}$ $0.261^{*}_{\pm 0.02}$ $0.462_{\pm 0.04}$ $0.620_{\pm 0.05}$
+SubQ Decomposition $035^{*}_{\pm 0.00} 0.0$$082_{\pm 0.03}$ $0.774^{*}_{\pm 0.08}$ $0.300_{\pm 0.05}$ $0.277^{*}_{\pm 0.01}$ $0.469_{\pm 0.04}$ $0.618_{\pm 0.05}$
+Claude-Haiku-4.5 (Anthropic, 2025a)
+Vanilla $040^{*}_{\pm 0.00} 0.0$$036_{\pm 0.02}$ $0.681^{*}_{\pm 0.11}$ $0.270_{\pm 0.04}$ $0.290^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.647_{\pm 0.05}$
+Chain-of-Thought $039^{*}_{\pm 0.00} 0.0$$036_{\pm 0.02}$ $0.748^{*}_{\pm 0.12}$ $0.260_{\pm 0.04}$ $0.284^{*}_{\pm 0.01}$ $0.669_{\pm 0.04}$ $0.608_{\pm 0.05}$
+Socratic Prompting $049^{*}_{\pm 0.00} 0.0$$091_{\pm 0.03}$ $0.671^{*}_{\pm 0.10}$ $0.270_{\pm 0.05}$ $0.303^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.324_{\pm 0.05}$
+SubQ Decomposition $043^{*}_{\pm 0.00} 0.0$$055_{\pm 0.02}$ $0.728^{*}_{\pm 0.11}$ $0.240_{\pm 0.04}$ $0.296^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.412_{\pm 0.05}$
+Gemini-2.5-Pro (team, 2025)
+Vanilla $0.097^{*}_{\pm 0.01}$ $0.227_{\pm 0.04}$ $0.901^{*}_{\pm 0.14}$ $0.280_{\pm 0.04}$ $0.216^{*}_{\pm 0.01}$ $0.408_{\pm 0.04}$ $0.598_{\pm 0.05}$
+Chain-of-Thought $0.101^{*}_{\pm 0.01}$ $0.209_{\pm 0.04}$ $0.971^{*}_{\pm 0.15}$ $0.270_{\pm 0.04}$ $0.219^{*}_{\pm 0.01}$ $0.392_{\pm 0.04}$ $0.588_{\pm 0.05}$
+Socratic Prompting $0.093^{*}_{\pm 0.01}$ $0.173_{\pm 0.04}$ $0.858^{*}_{\pm 0.15}$ $0.290_{\pm 0.05}$ $0.195^{*}_{\pm 0.02}$ $0.415_{\pm 0.04}$ $0.569_{\pm 0.05}$
+SubQ Decomposition $0.087^{*}_{\pm 0.01}$ $0.155_{\pm 0.04}$ $0.815^{*}_{\pm 0.15}$ $0.273_{\pm 0.04}$ $0.151^{*}_{\pm 0.01}$ $0.408_{\pm 0.04}$ $0.608_{\pm 0.05}$
+Gemini-2.5-Flash (team, 2025)
+Vanilla $081^{*}_{\pm 0.01} 0.0$$082_{\pm 0.03}$ $1.048^{*}_{\pm 0.18}$ $0.310_{\pm 0.05}$ $0.201^{*}_{\pm 0.01}$ $0.385_{\pm 0.04}$ $0.588_{\pm 0.05}$
+Chain-of-Thought $0.089^{*}_{\pm 0.01}$ $0.145_{\pm 0.03}$ $0.881^{*}_{\pm 0.15}$ $0.340_{\pm 0.05}$ $0.203^{*}_{\pm 0.01}$ $0.385_{\pm 0.04}$ $0.578_{\pm 0.05}$
+Socratic Prompting $084^{*}_{\pm 0.01} 0.0$$091_{\pm 0.03}$ $0.821^{*}_{\pm 0.14}$ $0.310_{\pm 0.05}$ $0.105^{*}_{\pm 0.01}$ $0.392_{\pm 0.04}$ $0.608_{\pm 0.05}$
+SubQ Decomposition $0.085^{*}_{\pm 0.01}$ $0.118_{\pm 0.03}$ $1.058^{*}_{\pm 0.17}$ $0.320_{\pm 0.05}$ $0.190^{*}_{\pm 0.01}$ $0.405_{\pm 0.04}$ $0.549_{\pm 0.05}$
+Table A3: Pipeline validation by domain. We report the mean accuracy for each stage of the pipeline and annotator agreement – Cohen’s $\kappa$.
 Domain $\mathcal{N}$ generated claims $\mathcal{N}$ aligned claims Claim Decomposition Accuracy Relevance Filtering Accuracy Expert Alignment Accuracy Cohen’s $\kappa$
@@
 Sepsis 108 66 0.900 0.852 0.894 0.3500
-Table A3: Pipeline validation by domain. We report the mean accuracy for each stage of the pipeline and annotator agreement – Cohen’s $\kappa$.
+Table A4: Model ablation study for evaluation pipeline.
+All experiments are run with vailla prompting.
+Across domains, model rankings are largely consistent across evaluators, indicating the robustness of our evaluation pipeline.
+Comparison of base models (explanation generator LLMs) across evaluator models (pipeline LLMs) for one dataset per domain.
+Best per evaluator is in bold; second-best is underlined.
+Evaluation Model Gemini-2.5-Flash-Lite GPT-5-Mini Qwen2.5-VL-7B-Instruct
+MassMaps
+GPT-5.2-Pro $0.749\pm 0.011$ $0.901\pm 0.008$ $0.653\pm 0.020$
+GPT-5-mini $0.644\pm 0.014$ $0.844\pm 0.011$ $0.570\pm 0.017$
+Claude-Opus-4.5 $0.558\pm 0.018$ $0.774\pm 0.017$ $0.468\pm 0.016$
+Claude-Haiku-4.5 $0.452\pm 0.018$ $0.742\pm 0.015$ $0.485\pm 0.017$
+Gemini-2.5-Pro $0.793\pm 0.013$ $0.801\pm 0.012$ $0.629\pm 0.017$
+Gemini-2.5-Flash $0.678\pm 0.014$ $0.836\pm 0.011$ $0.691\pm 0.017$
+Cholecystectomy
+GPT-5.2-Pro $0.681\pm 0.022$ $0.538\pm 0.030$ $0.686\pm 0.022$
+GPT-5-mini $0.777\pm 0.015$ $0.757\pm 0.016$ $0.643\pm 0.019$
+Claude-Opus-4.5 $0.665\pm 0.011$ $0.753\pm 0.011$ $0.666\pm 0.014$
+Claude-Haiku-4.5 $0.625\pm 0.012$ $0.677\pm 0.010$ $0.594\pm 0.012$
+Gemini-2.5-Pro $0.666\pm 0.012$ $0.697\pm 0.012$ $0.617\pm 0.014$
+Gemini-2.5-Flash $0.660\pm 0.012$ $0.685\pm 0.012$ $0.594\pm 0.012$
+Emotion
+GPT-5.2-Pro $0.729\pm 0.019$ $0.906\pm 0.012$ $0.728\pm 0.028$
+GPT-5-mini $0.730\pm 0.020$ $0.867\pm 0.019$ $0.751\pm 0.025$
+Claude-Opus-4.5 $0.796\pm 0.019$ $0.916\pm 0.016$ $0.752\pm 0.027$
+Claude-Haiku-4.5 $0.751\pm 0.024$ $0.881\pm 0.015$ $0.807\pm 0.022$
+Gemini-2.5-Pro $0.806\pm 0.023$ $0.888\pm 0.018$ $0.723\pm 0.029$
+Gemini-2.5-Flash $0.772\pm 0.026$ $0.900\pm 0.016$ $0.743\pm 0.028$
+Table A5: Expert-aligned claims (good and bad) across all T-FIX domains, with corresponding alignment scores and provided reasoning.
 Domain Claim Score (Category) Reasoning
@@
 [Bad] The patient’s lab results show an increased platelet count. 0.2 (SOFA Score Increase) SOFA score focuses on low platelet counts; increased count contradicts the criterion.
-Table A4: Expert-aligned claims (good and bad) across all T-FIX domains, with corresponding alignment scores and provided reasoning.
-Appendix A Extending T-FIX to a New Domain
-Though T-FIX covers a wide range of knowledge-intensive settings, it can easily be extended to additional domains.
-A key contribution of the T-FIX benchmark is the framework: we create a pipeline to score any free-form text explanation for expert alignment given a set of expert criteria.
-Additionally, we iterate extensively on all our prompt templates to ensure all T-FIX users need to do is input their task-specific details and perform no additional prompt engineering for good results.
-To add a new domain to T-FIX, we advise you to follow these steps:
-  1. 1.
-    Generate criteria: Use the deep research prompt template shown in Figure˜A4 to generate a list of expert alignment criteria for your domain.
-Optionally, have a domain expert vet the generated criteria.
-  2. 2.
-    Modify prompts: Modify the prompt templates outlined in Figure˜A1, Figure˜A2, and Figure˜A3 with your task description, few-shot examples, and generated expert criteria.
-  3. 3.
-    Run T-FIX: Plug in your prompts for each stage of the pipeline and run T-FIX on your dataset!
-We encourage you to contact the authors of this work if you need additional assistance setting up your custom domain.
-Appendix B Prompts for T-FIX Pipeline
-We show the prompts for Stage 1, 2, and 3 in Figure˜A1, Figure˜A2, and Figure˜A3, respectively.
+Appendix C Ablation for Evaluation Models
+To assess the robustness of our evaluation pipeline, we conduct ablations using different models as evaluators: closed-sourced GPT-5-mini (OpenAI, 2025b), Gemini-2.5-Flash-Lite (team, 2025), and an open-source model Qwen2.5-VL-7B-Instruct (Bai et al., 2025).
+We evaluate vanilla prompts on three datasets, each representing a different domain: Mass Maps (cosmology), Emotion (psychology), and Cholecystectomy (medicine).
+Table˜A4 summarizes the results:
+  • Mass Maps (cosmology): All three evaluators unanimously decide the explanations from two claude models, Claude-Opus-4.5 and Claude-Haiku-4.5, as the least aligned with expert intuition, while GPT and Gemini models align better.
+  • Cholecystectomy (medicine): All three models give low $\mathsf{T\textrm{-}FIXscore}$s to Claude-Haiku-4.5, Gemini-2.5-Pro, and Gemini-2.5-Flash.
+All three evaluation models give highest and second highest scores among two of GPT-5.2-Pro, GPT-5-Mini, and Claude-Opus-4.5.
+  • Emotion (psychology): All evaluators assign quite high scores across generation models, leading to minimal differences in rankings.
+On average, Claude-Opus-4.5 is consistently ranked as one of the top two models.
+Overall, these results indicate that our evaluation pipeline is robust to the choice of evaluation model.
+Appendix D Expert Validation
+D.1 Mass Maps
+Experts largely agreed with the model’s alignment on simple, descriptive claims that focused on observable visual patterns rather than interpretive reasoning.
+For instance, statements such as “The map displays a noticeable amount of blue and gray with significant red clustering” or mentions of “yellow spots” as dense regions were seen as well-aligned because they accurately reflected direct features in the image without overinterpreting them.
+The model was particularly consistent in correctly associating visible peaks and clusters with the relevant “Lensing Peak (Cluster) Abundance” category, showing strong performance on basic lensing-related observations.
+However, disagreements arose for more interpretive claims where the model either misclassified the category or underestimated alignment.
+For example, experts judged “The significant red clustering indicates areas with some dense structures” as completely aligned, while the model labeled it partial.
+Similarly, when the model associated “yellow spots” with fine-scale clumpiness instead of large cluster convergence, experts noted categorical confusion.
+At a higher level, experts appreciated the model’s handling of concrete, color- or peak-based features but found its phrasing vague when dealing with broader cosmological interpretations.
+They preferred precise terminology about structure type and scale, emphasizing that generic descriptions such as “noticeable fluctuations” failed to convey sufficient scientific specificity.
+D.2 Supernova
+Experts agreed with alignment scores for claims describing basic observable behaviors, such as “The light curve is followed by a gradual decline” or “The multi-wavelength observation shows a subsequent decline across all wavelengths,” which correctly corresponded to the “Monotonic Flux Trend” category.
+These statements were straightforward and descriptive, and the system’s partial alignment assessments matched expert expectations.
+Disagreement arose for claims that overinterpreted observational patterns.
+For example, “The decline across all wavelengths corroborates the classification as a Type Ia supernova” was given full alignment by the model, but experts argued it should be partial, as such behavior is characteristic of transient events generally, not specific to Type Ia supernovae.
+Experts praised the model’s ability to identify key distinguishing features, like “The light curve shows a rapid rise to a peak,” which effectively differentiates supernovae from other celestial phenomena.
+However, they were less satisfied with claims that only ruled out other classes (e.g., RR Lyrae, AGN, Mira) without affirmatively justifying the target classification.
+They stressed that negative reasoning — describing what an instance is not — does not equate to correctly identifying what it is.
+D.3 Politeness
+Experts generally agreed with the model’s alignment on explicit and lexical politeness cues.
+For example, both the model and experts assigned full alignment to claims like “The speaker expresses appreciation by saying ‘thank you,’” and “The utterance softens the request through the word ‘please.’” These cases show that the model reliably recognizes direct politeness indicators that have stable, context-independent meanings.
+Disagreement arose for subtler forms of politeness that rely on social context or pragmatic interpretation.
+Experts noted that the model often overestimated alignment for hedges and modal constructions, such as I think, maybe, or could you, which did not always convey genuine politeness signal.
+For instance, the model rated “The use of ‘we’ conveys solidarity” as fully aligned, while experts marked it partial because the “we” referred to an institution rather than shared identity.
+At a higher level, experts appreciated the model’s consistency and strong performance on overt politeness expressions but emphasized its limited pragmatic sensitivity.
+Overall, experts found the LLM trustworthy for identifying surface politeness cues but less dependable when interpreting indirect, contextual, or culturally grounded politeness strategies.
+D.4 Emotion
+Experts generally agreed with the model’s alignment for surface-level emotional cues.
+Claims referencing Emotion Words, Emojis, Expressive Punctuation, Humor Markers were accurately scored by both the model and experts.
+The model also showed strong reliability in labeling claims related to Gratitude and Praise.
+These results suggest that the model can be trusted to judge alignment accurately when the emotion is lexically or visually explicit.
+Disagreement emerged in more nuanced emotional expressions where context determined the emotional interpretation.
+Experts noted that the model often overestimated alignment for Valence and Arousal categories, particularly when phrasing was neutral or ambiguous.
+Claims involving subtle or mixed emotions, such as those involving Relief or Affection, were rarely labeled partial, as the model treated any positive cue as a strong emotional expression.
+At a broader level, experts were happy with the model’s precision in labeling but noted limited sensitivity to gradation and context.
+The system performed best when emotions were unambiguous and directly expressed, but faltered when affect was implied, ironic, or intertwined with multiple sentiments.
+D.5 Laparoscopic Cholecystectomy
+Experts generally agreed with the model when it judged purely positional claims as not aligned with expert reasoning.
+They emphasized that referring only to a portion of the screen does not constitute surgically meaningful evidence for safety or risk.
+For example, the claim “The unsafe region occupies the lower to center-right portion of the image” was considered uninformative, as it describes only screen location rather than anatomically or surgically relevant cues.
+In such cases, experts agreed that alignment scores correctly reflected the lack of expert-valid reasoning.
+Disagreement arose when the model assigned alignment to an incorrect anatomical category or relied on implausible visual inferences.
+For instance, experts judged the statement “The lower to center-right portion of the image shows darker, more vascular tissue” as less related to the “Peritoneal plane” category and more indicative of possible porta hepatis or vascular structures—although they emphasized that these structures cannot be directly identified from the image.
+Experts also noted that some unsafe-zone criteria, such as using the cystic lymph node as a landmark, are not visible in all cases and therefore difficult to apply consistently.
+In several instances, they pointed out that the model hallucinated tissues or structures that were not actually present in the image.
+At a higher level, experts appreciated that the model correctly rejected claims based solely on screen position as misaligned with surgical intuition.
+However, they found the model less reliable when reasoning about anatomical structures whose visibility or relevance was uncertain.
+They stressed that many key anatomical landmarks cannot be directly observed in laparoscopic views and that some structures are far more informative than others for safety assessment.
+In particular, they highlighted the cystic lymph node and the estimation of the gallbladder infundibulum as useful landmarks, while cautioning against overinterpreting poorly visible or anatomically ambiguous regions.
+D.6 Cardiac
+Experts found most alignment scores accurate and consistent with their own judgments, highlighting the model’s general reliability in this domain.
+They appreciated the structural clarity and systematic nature of the pipeline, noting that alignment scores often captured correct relationships between claims and expert categories.
+Nevertheless, they pointed out that while scores were often correct, the model’s reasoning behind them was sometimes incomplete or slightly off, revealing a disconnect between the correctness of the label and the explanatory rationale.
+At a broader level, experts valued the coherence of the approach and the alignment consistency across claims.
+However, they criticized redundancy and excessive complexity in some claims.
+For example, two statements about ECG rhythm regularity were nearly identical, suggesting overgeneration.
+They also noted that some claims were too dense and should be broken into smaller, more specific parts.
+Additional concerns included inaccuracies in certain category definitions, such as labeling 30 years old as “advanced age”, and insufficient background context for some explanations.
+Overall, they viewed the system as strong in structure but in need of refinement in content granularity and contextual accuracy.
+D.7 Sepsis
+Experts agreed with most alignment scores, particularly those labeled partial, since the claims captured correct but incomplete aspects of the clinical reasoning process.
+For instance, “One risk factor for sepsis is advanced age” was appropriately labeled as partial under “Elderly Susceptibility,” as it was factually true but lacked the specific clinical threshold (age $\geq$ 65).
+Similarly, claims about single vital sign abnormalities were correctly labeled as partial since sepsis criteria like SIRS require multiple indicators.
+However, experts disagreed when the model evaluated isolated claims that should have been considered in combination.
+Statements like “A high triage temperature indicates fever” and “An elevated respiratory rate is another risk factor” were accurate but insufficient alone to establish “SIRS Positivity,” leading to misalignment between formal criteria and contextual reasoning.
+Additionally, experts highlighted that the LLM would generate explanations that involved risk factors that are not the same as diagnostic criteria for sepsis — while risk factors increase the likelihood of sepsis, there is no causal relationship.
+At a higher level, experts appreciated the clarity of claims tied to objective measures such as qSOFA scores and age but highlighted a persistent tension between rule-based alignment and real-world clinical reasoning.
+They emphasized that diagnostic logic in practice is more flexible than rigid scoring systems, exposing a conceptual gap that the model did not fully capture.
+Appendix E Prompts for T-FIX Pipeline
+We show the prompts for Stage 1, 2, and 3 in Figure˜A1, Figure˜A2, Figure˜A3, and Figure˜A4, respectively.
 These prompts show a high-level template that was used by all domains.
 In practice, authors iterated multiple times on each domain’s prompts, experimenting with the instruction wording and few-shot examples that yielded the best possible results.
-Appendix C Ablation for Evaluation Models
-To assess the robustness of our evaluation pipeline, we conduct ablations using different models as evaluators: closed-sourced GPT-4o (openai2024gpt4ocard), Gemini-2.0-Flash (google2024gemini2flash), and an open-source model Qwen2.5-VL-7B-Instruct (Qwen2.5-VL).
-We evaluate vanilla prompts on three datasets, each representing a different domain: Mass Maps (cosmology), Emotion (psychology), and Cholecystectomy (medicine).
-Table˜A1 summarizes the results:
-  • Mass Maps (cosmology): All three evaluators unanimously rank Claude-3.5-Sonnet as the best model and o1 as the second best.
-  • Cholecystectomy (medicine): GPT-4o and Qwen2.5-VL-7B-Instruct both rank o1 as the best and Qwen2.5-VL-7B-Instruct as the second best.
-Gemini-2.0-Flash is more conservative, labeling all claims as irrelevant and assigning all models a score of 0.
-  • Emotion (psychology): All evaluators assign similar scores across generation models, leading to minimal differences in rankings.
-Overall, these results indicate that our evaluation pipeline is robust to the choice of evaluation model.
+Appendix F Framework Development Process
+In this section, we detail how we engineered all the prompts we tried for each of the 4 prompts in our T-FIX pipeline, what worked and what did not work, and how expert feedback was incorporated throughout the process.
+F.1 Atomic claim extraction
+The atomic claim extraction step decomposes each model-generated explanation into a set of atomic, independently interpretable claims for downstream relevance filtering and expert alignment.
+We designed structured prompts for each domain that enforce atomicity, standalone interpretability, and faithfulness to the original explanation.
+What did not work. Given we began with the refined atomic claim extraction prompts from Wanner et al. (2024), this step needed little prompt engineering to work as expected.
+What worked.
+We found that explicitly defining atomicity, standalone interpretation, and faithfulness within the prompt substantially improved the quality and consistency of extracted claims.
+Providing concrete in-context examples of correctly decomposed explanations for each domain further stabilized output structure and reduced stylistic variability across runs.
+Additionally, instructing the model to output only a list of claims with no additional commentary simplified downstream processing and facilitated automated relevance filtering and alignment.
+F.2 Relevancy filtering
+In the relevancy filtering step, we determine whether each extracted claim should be retained as supporting evidence for the model’s prediction.
+We prompted the LLM to label a claim as relevant only if it (i) is supported by the provided input and (ii) directly contributes to explaining the predicted outcome.
+This formulation filters out unsupported statements, generic background knowledge, and tautological claims that restate the prediction without adding evidence, yielding a cleaner set of claims for downstream grouping and expert alignment.
+What did not work.
+Prompts that framed relevance in vague terms (e.g., “important,” “useful,” or “related”) led to inconsistent decisions and frequent retention of statements that were not meaningful for the explanation — restating the prediction, describing the input, describing the task, etc.
+We required multiple rounds of prompt iterations and a clear definition of relevance to get this stage to work with high accuracy and reliability.
+What worked.
+We found that defining relevance with two explicit criteria, (i) the claim must be supported by the input and (ii) it must help explain the model’s prediction, yielded the most consistent filtering behavior.
+Requiring the model to generate a a reasoning to support its decision improved performance and made failure modes easier to diagnose during expert review.
+Providing a set of diverse examples further helped the model decide which claims were speculative, unsupported by the input, or not meaningful for the explanation.
+F.3 Step 3: Claim grouping
+This step required the highest number of iterations, as it is where we most incorporated domain expert feedback.
+We have broken down our process to the main components of prompt generation for expert knowledge.
+Identifying Expert Categories.
+First, we used o3 deep research (OpenAI, 2025c) to seed our list of expert categories for each domain.
+We verified the model was pulling from reputable, existing literature, and manually ensured all categories made semantic sense and covered the areas that the experts expected.
+Following this, there were multiple (at least 2 per domain) iterations of review with the domain experts themselves for the determination of what constitutes the “ideal” expert categories for each domain.
+Working with the experts, we facilitated iterations where categories were modified, removed, and replaced, until experts reached consensus and the categories were finalized.
+What did not work.
+Our initial setup of this pipeline did not involve the claim grouping step at all; we evaluated alignment of individual claims and aggregated the results.
+However, our domain experts pointed out that expert alignment is often determined by the joint content of multiple claims: claims can depend on one another, and some criteria are inherently compositional.
+For example, SIRS positivity in sepsis diagnosis requires meeting at least two symptoms from a predefined list, which may be expressed across several separate claims.
+Incorporating this feedback, we refactored our pipeline to incorporate claim grouping before expert alignment.
+This way, all claims related to a category could be judged together.
+Initial prompts returned groupings that were too selective; claims rarely belonged to multiple categories and many categories had zero related claims.
+Without explicit instructions to consider all claims before selecting a subset, GPT-5-mini sometimes defaulted to selecting only the most obviously relates claims and ignored supporting, but less obvious evidence.
+What worked.
+After finalizing the expert criteria, we developed prompts that allowed us to group all relevant, filtered claims into criteria-level groups.
+We allowed the claims to belong to multiple criteria, as certain facts or statements in our explanations were referenced by multiple criteria.
+What worked well was curating a diverse set of examples with varying downstream groupings: most claims in a category, around 1/2 of the claims in a category, and no claims in a category.
+GPT-5-mini erred on the side of fewer claims in a category, so we included more examples of categories with large claim groupings to account for this.
+Upon encouraging GPT-5-mini to be more generous and inclusive with the grouping, we were satisfied with the results.
+F.4 Expert alignment of claim groups
+This step also required a high amount of iteration, and domain expert feedback was heavily used in creating our final alignment prompts.
+What did not work.
+The first attempt of our claim to expert category alignment mapping was straightforward and simple.
+For each claim, we assigned it to the category that it most aligned with according to the LLM, and then calculated alignment.
+If the claim was completely not aligned with any of the expert categories, then it was omitted.
+We originally used a 0 to 1 scale with increments of 0.1 to score how aligned a claim was to a given expert category, where 1 is completely aligned and 0 is not aligned at all.
+This turned out to be too fine-grained a measure because it was unclear and ambiguous how to determine the difference between claims that got scores of 0.4 and 0.6, for example Moreover, numeric scores encouraged spurious precision and made it difficult for domain experts to interpret or validate the assignments.
+We also found that even with explicit definitions of the scale, the LLM’s scoring was unstable across runs and sensitive to minor prompt changes.
+What worked.
+Based on feedback from our experts, we added the claim grouping step (see above).
+This group-based formulation better captured how expert reasoning often relies on multiple supporting statements rather than isolated claims.
+After experimentation, we also determined that using discrete alignment labels: none, partial, and complete alignments for each claim group was the most effective approach, because it provided clear semantic interpretation of how well a claim reflected an expert category.
+This formulation reduced ambiguity and improved consistency compared to numeric scoring schemes.
+Examples of claim groups with partial alignment were especially important in helping GPT-5-mini handle borderline cases where claims were related to a category but lacked sufficient specificity.
+Similar to previous steps, requiring the LLM to provide a short justification for each alignment decision helped debug errors and improved performance.
+What did not work.
+The initial versions of this prompt frequently resulted in the inclusion of too many claims per category, including irrelevant or weakly related statements.
+The LLM also tended to omit subtle but important claims when they were phrased indirectly or lacked explicit category keywords.
+Additionally, without explicit instructions to consider all claims before selecting a subset, the model sometimes defaulted to selecting only the most obvious claims and ignored secondary evidence.
+This required prompt revisions that enforced exhaustive consideration of the claim set and stricter inclusion criteria.
 Prompt
@@
 Prompt
-Figure A2: Prompt Template for Stage 2: Relevancy Filtering
+Figure A2: Prompt Template for Stage 2A: Relevancy Filtering
+Prompt
+Figure A3: Prompt Template for Stage 2B: Claim Grouping
 Prompt
@@
 Figure A6: Baseline Prompting Strategies.
-Appendix D T-FIX Datasets: Additional Details
+Appendix G T-FIX Datasets: Additional Details
 Prompt
 Figure A7: MassMaps Explanation Prompt
-D.1 Mass Maps
-Task.
+G.1 Mass Maps
 […]
@@
 We randomly sampled 100 examples from the MassMaps test set.
-To ensure compatibility with LLMs like GPT-4o, which operate on a 32×32 patch size, we upsampled each image by a factor of 11 to preserve spatial detail and avoid patch-level compression.
+To ensure compatibility with LLMs like GPT-5-mini, which operate on a 32×32 patch size, we upsampled each image by a factor of 11 to preserve spatial detail and avoid patch-level compression.
 Instead of raw pixel values, we applied a colormap based on expert-defined intensity thresholds used to identify key cosmological features such as voids and clusters.
@@
 Figure A7 shows the prompt used to generate LLM explanations for predicting $\Omega_{m}$ and $\sigma_{8}$.
-We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
+We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
 The prompt includes a description of how pixel values are mapped to colors, as well as the valid ranges for $\Omega_{m}$ and $\sigma_{8}$.
 Without this range, models tend to default to common values (e.g., 0.3 for $\Omega_{m}$, 0.8 for $\sigma_{8}$), reducing response variability.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
@@
 Figure A8: Supernova Explanation Prompt
-D.2 Supernova
+G.2 Supernova
 Task.
 The objective is to classify astrophysical objects using time-series data comprising observation times (Modified Julian Dates), wavelengths (filters), flux values, and corresponding flux uncertainties.
-We use data from the PLAsTiCC challenge (theplasticcteam2018photometric), where the model must predict one of 14 astrophysical classes.
+We use data from the PLAsTiCC challenge (Team et al., 2018), where the model must predict one of 14 astrophysical classes.
 Data Selection & Preprocessing.
@@
 Figure A8 shows the prompt used to generate explanations for classifying astronomical objects.
-We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
+We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
 The prompt includes a description of the input plot as a multivariate time series and provides the full list of possible class labels to guide the model’s predictions.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
@@
     Monotonic flux trends: Locally smooth, monotonic flux trends across one or multiple bands (plateaus, linear decays) capture physical evolution stages and help distinguish SNII‑P, SNII‑L, and related classes.
-D.3 Politeness
+G.3 Politeness
 Task.
 Understanding how linguistic styles, like politeness, vary across cultures is necessary for building better communication, translation, and conversation-focused systems.
-holmes2012politeness; havaldar2023multilingual.
-Today’s LLMs exhibit large amounts of cultural bias havaldar2024building, and understanding nuances in cultural differences can help encourage cultural adaptation in models.
-We use the holistic politeness dataset from havaldar2025comparingstyleslanguagescrosscultural, which consists of conversational utterances between editors from Wikipedia talk pages, annotated by native speakers from four distinct cultures.
+(Holmes, 2012; Havaldar et al., 2023b).
+Today’s LLMs exhibit large amounts of cultural bias (Havaldar et al., 2024), and understanding nuances in cultural differences can help encourage cultural adaptation in models.
+We use the holistic politeness dataset from Havaldar et al.
+(2023a), which consists of conversational utterances between editors from Wikipedia talk pages, annotated by native speakers from four distinct cultures.
 Data Selection & Preprocessing.
@@
 Figure A9: Politeness Explanation Prompt
-Explanation Prompt.
-We show the prompt in Figure˜A8. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Explanation Prompt. We show the prompt in Figure˜A9. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
@@
     Ingroup Language and Informality: Use of group-identifying slang or casual expressions (“mate,” “dude,” “bro”) may foster solidarity or seem disrespectful, depending on relational norms.
-D.4 Emotion
+G.4 Emotion
 Task.
 Understanding and classifying emotion is important for tasks like therapy, mental health diagnoses, etc.
-denzin1984understanding.
-Emotion is often expressed implicitly, and understanding such cues can also aid in building LLM systems that handle implied language understanding well havaldar2025entailed.
-We use the GoEmotions dataset from demszky2020goemotions, consisting of Reddit comments that have been human-annotated for one of 27 emotions (or neutral, if no emotion is present).
+(Denzin, 1984).
+Emotion is often expressed implicitly, and understanding such cues can also aid in building LLM systems that handle implied language understanding well (Havaldar et al., 2025).
+We use the GoEmotions dataset from Demszky et al.
+(2020), consisting of Reddit comments that have been human-annotated for one of 27 emotions (or neutral, if no emotion is present).
 Data Selection & Preprocessing.
@@
 Figure A10: Emotion Explanation Prompt
-Explanation Prompt.
-We show the prompt in Figure˜A9. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Explanation Prompt. We show the prompt in Figure˜A10. We replace ‘‘[BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
@@
     Relief Indicators: Release phrases like “phew,” “finally over,” or “what a relief” mark relief after stress ends.
-D.5 Laparoscopic Cholecystectomy Surgery.
+G.5 Laparoscopic Cholecystectomy
 Task.
 The task is to identify the safe and unsafe regions for incision.
-We used the open-source subset of data from (madani2022artificial), which consists of surgeon-annotated images taken from video frames from the M2CAI16 workflow challenge (stauder2016tum) and Cholec80 (twinanda2016endonet) datasets.
+We used the open-source subset of data from (Madani et al., 2022), which consists of surgeon-annotated images taken from video frames from the M2CAI16 workflow challenge (Stauder et al., 2016) and Cholec80 (Twinanda et al., 2016) datasets.
 This consists of 1015 surgeon-annotated images.
@@
 The LLM was to then output two lists with numbers from this range: a “safe list” that denotes its prediction of the safe region, and an “unsafe list” predicting the unsafe region.
-Explanation Prompt.
-We show the prompt in Figure˜A10. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Explanation Prompt. We show the prompt in Figure˜A11. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
-    Calot’s triangle cleared - Hepatocystic triangle must be fully cleared of fat/fibrosis so that its boundaries are unmistakable.
+    Calot’s triangle cleared: Hepatocystic triangle must be fully cleared of fat/fibrosis so that its boundaries are unmistakable.
   2. 2.
-    Cystic plate exposed - The lower third of the gallbladder must be dissected off the liver to reveal the shiny cystic plate and ensure the correct dissection plane.
+    Cystic plate exposed: The lower third of the gallbladder must be dissected off the liver to reveal the shiny cystic plate and ensure the correct dissection plane.
   3. 3.
-    Only two structures visible - Only the cystic duct and cystic artery should be seen entering the gallbladder before any clipping or cutting.
+    Only two structures visible: Only the cystic duct and cystic artery should be seen entering the gallbladder before any clipping or cutting.
   4. 4.
-    Above the R4U line - Dissection must remain cephalad to an imaginary line from Rouviere’s sulcus to liver segment IV to avoid the common bile duct.
+    Above the R4U line: Dissection must remain cephalad to an imaginary line from Rouviere’s sulcus to liver segment IV umbilical fissure to avoid the common bile duct.
+Dissection should be carried out along the inferior edge of the gallbladder (well above the line of safety).
   5. 5.
-    Safe distance from common bile duct - There should be sufficient distance between the common bile duct and the gallbladder wall to ensure safe dissection.
+    Infundibulum start point: Dissection can begin at the gallbladder infundibulum-cystic duct junction to stay in safe tissue planes, or at the lateral or medial edges of the gallbladder above Rouviere’s sulcus or along the cystic plate to get mobility of the gallbladder first.
   6. 6.
-    Infundibulum start point - Dissection should begin at the gallbladder infundibulum-cystic duct junction to stay in safe tissue planes.
+    Peritoneal plane: When separating the gallbladder from the liver, stay in the avascular peritoneal cleavage plane.
   7. 7.
-    Subserosal plane stay - When separating the gallbladder from the liver, stay in the avascular subserosal cleavage plane under the serosal fat layer.
+    Cystic lymph node (calot’s node) guide: Identify the cystic lymph node and clip the artery on the gallbladder side of the node to avoid injuring the hepatic artery.
   8. 8.
-    Cystic lymph node guide - Identify the cystic lymph node and clip the artery on the gallbladder side of the node to avoid injuring the hepatic artery.
+    No division without ID: Never divide any duct or vessel until it is unequivocally identified as the cystic structure entering the gallbladder.
   9. 9.
-    No division without ID - Never divide any duct or vessel until it is unequivocally identified as the cystic structure entering the gallbladder.
-  10. 10.
-    Inflammation bailout - If dense scarring or distorted anatomy obscures Calot’s triangle, convert to a subtotal "fundus-first" approach rather than blind cutting.
-  11. 11.
-    Aberrant artery caution - Preserve any large or tortuous artery (e.g., a Moynihan’s hump) that might be mistaken for the cystic artery.
+    Inflammation bailout: If dense scarring or distorted anatomy obscures Calot’s triangle, convert to open surgery or a fenestrated subtotal approach rather than blind cutting.
 Prompt
@@
 Figure A12: Cardiac Explanation Prompt
-D.6 Cardiac Arrest
+G.6 Cardiac Arrest
 Task.
 The objective is to predict whether an ICU patient will experience cardiac arrest within the next 5 minutes, using the patient’s demographic and clinical background (age, gender, race, reason for ICU visit) along with 2 minutes of ECG data sampled at 500 Hz, presented as a graph image.
-This framing aligns with cardiology literature, which suggests that short ECG windows (30 seconds to a few minutes) are sufficient for reliable prediction nussinovitch2011reliability.
+This framing aligns with cardiology literature, which suggests that short ECG windows (30 seconds to a few minutes) are sufficient for reliable prediction (Nussinovitch et al., 2011).
 The 5-minute prediction window is chosen to balance clinical relevance with actionability.
 Data Selection & Preprocessing.
-We use ECG and visit data from the open-source Multimodal Clinical Monitoring in the Emergency Department (MC-MED) Dataset kansal2025mcmed.
+We use ECG and visit data from the open-source Multimodal Clinical Monitoring in the Emergency Department (MC-MED) Dataset (Kansal et al., 2025).
 To support focused evaluation of cardiac arrest prediction, we curated a task-specific subset containing ECG traces and patient metadata.
 The data curation pipeline proceeded as follows. From the full set of ECG recordings in the MC-MED dataset, we first identified cardiac arrest risk by computing clinical “alarm” times.
-Prior work shows that vital sign abnormalities are predictive of outcomes Candel2022-bo; chen2023multimodal.
+Prior work shows that vital sign abnormalities are predictive of outcomes (Candel et al., 2022; Chen et al., 2023).
 We defined an alarm at any timestamp where three or more of the following vital signs were outside normal range within a two-minute window—a condition known clinically as decompensation:
@@
 Figure A12 shows the prompt used to generate explanations for predicting whether an ICU patient will experience cardiac arrest within 5 minutes, based on 2 minutes of ECG data along with age, gender, race, and ICU admission reason.
-We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
+We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
 The ECG is provided as a graph image of p-signal values sampled at 500 Hz over a 2-minute window, with labeled axes.
 While we considered supplying the raw signal as text, the input token limits of current LLMs made this infeasible.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
-    Ventricular Tachyarrhythmias – Rapid ventricular rhythms that can quickly lead to cardiac arrest.
+    Ventricular Ectopy / NSVT: Runs of non-sustained ventricular tachycardia (NSVT) or frequent premature ventricular contractions may indicate electrical instability in critically ill patients, particularly those with underlying coronary disease or cardiomyopathy.
+While NSVT is generally considered a more benign form of ventricular tachyarrhythmia, its presence reflects transient abnormal rhythms originating from the lower chambers and does not necessarily always signal impending cardiac arrest.
   2. 2.
-    Ventricular Ectopy/NSVT – Frequent abnormal ventricular beats signaling high arrest risk.
+    Bradycardia or Heart-Rate Drop: The onset of significant bradycardia or a sudden $\geq 30\%$ decline in heart rate is a well-documented precursor to in-hospital cardiac arrest (often preceding pulseless electrical activity or asystole) and should be treated as an alarm sign.
   3. 3.
-    Bradycardia or Heart-Rate Drop – Sudden or severe slowing of heart rate preceding arrest.
+    QRS Widening (Conduction Delay): New or progressive prolongation of the QRS duration on the ECG is an ominous finding in the ICU, often observed in the minutes before cardiac arrest and associated with higher mortality due to deteriorating ventricular conduction.
   4. 4.
-    Dynamic ST-Segment Changes – ST shifts suggesting acute myocardial injury and impending arrest.
+    Dynamic ST-Segment Changes: Acute ischemic changes on continuous ECG (notably ST-segment elevation or depression) signal low blood flow in the coronary arteries, indicating myocardial infarction or injury, and may precede imminent ventricular fibrillation or cardiac arrest.
+As blood flow continues to decrease, the ischemic heart can fibrillate or go into VT/VF (i.e., cardiac arrest).
+Although ST segment changes are common, the link to cardiac arrest is rare but possible.
   5. 5.
-    Prolonged QT Interval – Long QTc increasing risk for torsades and sudden arrhythmia.
+    QTc Prolongation: Prolongation of the corrected QT (QTc) interval reflects abnormal ventricular repolarization and increases vulnerability to malignant arrhythmias such as torsades de pointes.
+QTc prolongation may result from medications, electrolyte disturbances (e.g., hypokalemia or hypomagnesemia), or underlying cardiac disease and is a known precursor to sudden cardiac arrest.
   6. 6.
-    Severe Hyperkalemia Signs – ECG changes from high potassium predicting arrest, especially among patients on dialysis / end stage renal disease.
+    Severe Hyperkalemia Signs: Electrocardiographic signs of severe hyperkalemia (such as peaked T-waves, loss of P-waves, and a widening QRS complex) herald an impending arrest – as potassium levels rise, the ECG may evolve to a sine-wave pattern and typically culminate in ventricular fibrillation or asystole without immediate intervention.
+Hyperkalemia is a frequent cause of in-hospital cardiac arrest especially among patients on dialysis / end stage renal disease.
+Looking for signs of hyperkalemia can be important to understand risk of cardiac arrest, especially in selected populations.
   7. 7.
-    Advanced Age – Older age strongly correlates with higher arrest likelihood.
+    Advanced Age: Increasing age is a major risk factor for cardiac arrest, with events being uncommon in younger patients and substantially more frequent in older ICU populations.
+Patients over 65 years are more prone to sudden deterioration due to reduced physiologic reserve and a higher burden of cardiovascular disease.
   8. 8.
-    Male Sex – Males have a higher overall risk of cardiac arrest.
+    Male Sex: Male gender is associated with a higher incidence of cardiac arrest, as most cardiac arrests occur in men (with women’s risk rising post-menopause).
   9. 9.
-    Underlying Cardiac Disease – Preexisting heart disease increases arrest susceptibility.
+    Underlying Cardiac Disease: The presence of serious cardiac conditions – such as coronary artery disease (especially a recent myocardial infarction) or severe heart failure – greatly elevates short-term cardiac arrest risk by creating an electrically and hemodynamically unstable myocardium.
   10. 10.
-    Critical Illness (Sepsis/Shock) – Severe infections or shock states elevate arrest risk through systemic instability.
+    Critical Illness (Sepsis/Shock): Severe sepsis or septic shock substantially raises the likelihood of cardiac arrest in the near term by causing hypoxia, hypotension, and metabolic derangements that often lead to pulseless electrical activity or asystole.
 Prompt
 Figure A13: Sepsis Explanation Prompt
-D.7 Sepsis
+G.7 Sepsis
 Task.
@@
 Each input is a time series of records containing a timestamp, the name of a physiological measurement or medication, and its value.
-Data Selection & Preprocessing.
-We used data from the publicly available MC-MED dataset kansal2025mcmed and curated a task-specific subset for sepsis prediction.
+Data Selection & Preprocessing. We used data from the publicly available MC-MED dataset (Kansal et al., 2025) and curated a task-specific subset for sepsis prediction.
 To label a patient as high risk for sepsis, we followed standard clinical definitions requiring three conditions: (1) evidence of infection, indicated by either a blood culture being drawn or at least two hours of antibiotic administration; (2) signs of organ dysfunction, defined by a SOFA score $\geq$2 within 48 hours of suspected infection, based on abnormalities in respiratory, coagulation, liver, cardiovascular, neurological, or renal function; and (3) presence of fever, with a recorded temperature $\geq$38.0°C (100.4°F).
@@
 Figure A13 shows the prompt used to generate LLM explanations for sepsis risk prediction.
-We substitute [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5.
+We substitute [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6.
 The prompt includes a description of the EHR input format: each time-series record consists of a timestamp, a measurement or medication name, and its value.
-Expert Criteria.
-The expert-validated criteria for expert alignment calculation are listed below:
+Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
   1. 1.
@@
     Early Antibiotic/Culture Orders (within 2 hours): Administration of broad-spectrum antibiotics or drawing of blood cultures within the first 2 hours signifies clinician suspicion of serious infection and should anchor sepsis risk assessment.
-Appendix E Expert Validation
-E.1 Mass Maps
-Experts largely agreed with the model’s alignment on simple, descriptive claims that focused on observable visual patterns rather than interpretive reasoning.
-For instance, statements such as “The map displays a noticeable amount of blue and gray with significant red clustering” or mentions of “yellow spots” as dense regions were seen as well-aligned because they accurately reflected direct features in the image without overinterpreting them.
-The model was particularly consistent in correctly associating visible peaks and clusters with the relevant “Lensing Peak (Cluster) Abundance” category, showing strong performance on basic lensing-related observations.
-However, disagreements arose for more interpretive claims where the model either misclassified the category or underestimated alignment.
-For example, experts judged “The significant red clustering indicates areas with some dense structures” as completely aligned, while the model labeled it partial.
-Similarly, when the model associated “yellow spots” with fine-scale clumpiness instead of large cluster convergence, experts noted categorical confusion.
-At a higher level, experts appreciated the model’s handling of concrete, color- or peak-based features but found its phrasing vague when dealing with broader cosmological interpretations.
-They preferred precise terminology about structure type and scale, emphasizing that generic descriptions such as “noticeable fluctuations” failed to convey sufficient scientific specificity.
-E.2 Supernova
-Experts agreed with alignment scores for claims describing basic observable behaviors, such as “The light curve is followed by a gradual decline” or “The multi-wavelength observation shows a subsequent decline across all wavelengths,” which correctly corresponded to the “Monotonic Flux Trend” category.
-These statements were straightforward and descriptive, and the system’s partial alignment assessments matched expert expectations.
-Disagreement arose for claims that overinterpreted observational patterns.
-For example, “The decline across all wavelengths corroborates the classification as a Type Ia supernova” was given full alignment by the model, but experts argued it should be partial, as such behavior is characteristic of transient events generally, not specific to Type Ia supernovae.
-Experts praised the model’s ability to identify key distinguishing features, like “The light curve shows a rapid rise to a peak,” which effectively differentiates supernovae from other celestial phenomena.
-However, they were less satisfied with claims that only ruled out other classes (e.g., RR Lyrae, AGN, Mira) without affirmatively justifying the target classification.
-They stressed that negative reasoning — describing what an instance is not — does not equate to correctly identifying what it is.
-E.3 Politeness
-Experts generally agreed with the model’s alignment on explicit and lexical politeness cues.
-For example, both the model and experts assigned full alignment to claims like “The speaker expresses appreciation by saying ‘thank you,’” and “The utterance softens the request through the word ‘please.’” These cases show that the model reliably recognizes direct politeness indicators that have stable, context-independent meanings.
-Disagreement arose for subtler forms of politeness that rely on social context or pragmatic interpretation.
-Experts noted that the model often overestimated alignment for hedges and modal constructions, such as I think, maybe, or could you, which did not always convey genuine politeness signal.
-For instance, the model rated “The use of ‘we’ conveys solidarity” as fully aligned, while experts marked it partial because the “we” referred to an institution rather than shared identity.
-At a higher level, experts appreciated the model’s consistency and strong performance on overt politeness expressions but emphasized its limited pragmatic sensitivity.
-Overall, experts found the LLM trustworthy for identifying surface politeness cues but less dependable when interpreting indirect, contextual, or culturally grounded politeness strategies.
-E.4 Emotion
-Experts generally agreed with the model’s alignment for surface-level emotional cues.
-Claims referencing Emotion Words, Emojis, Expressive Punctuation, Humor Markers were accurately scored by both the model and experts.
-The model also showed strong reliability in labeling claims related to Gratitude and Praise.
-These results suggest that the model can be trusted to judge alignment accurately when the emotion is lexically or visually explicit.
-Disagreement emerged in more nuanced emotional expressions where context determined the emotional interpretation.
-Experts noted that the model often overestimated alignment for Valence and Arousal categories, particularly when phrasing was neutral or ambiguous.
-Claims involving subtle or mixed emotions, such as those involving Relief or Affection, were rarely labeled partial, as the model treated any positive cue as a strong emotional expression.
-At a broader level, experts were happy with the model’s precision in labeling but noted limited sensitivity to gradation and context.
-The system performed best when emotions were unambiguous and directly expressed, but faltered when affect was implied, ironic, or intertwined with multiple sentiments.
-E.5 Cardiac
-Experts found most alignment scores accurate and consistent with their own judgments, highlighting the model’s general reliability in this domain.
-They appreciated the structural clarity and systematic nature of the pipeline, noting that alignment scores often captured correct relationships between claims and expert categories.
-Nevertheless, they pointed out that while scores were often correct, the model’s reasoning behind them was sometimes incomplete or slightly off, revealing a disconnect between the correctness of the label and the explanatory rationale.
-At a broader level, experts valued the coherence of the approach and the alignment consistency across claims.
-However, they criticized redundancy and excessive complexity in some claims.
-For example, two statements about ECG rhythm regularity were nearly identical, suggesting overgeneration.
-They also noted that some claims were too dense and should be broken into smaller, more specific parts.
-Additional concerns included inaccuracies in certain category definitions, such as labeling 30 years old as “advanced age”, and insufficient background context for some explanations.
-Overall, they viewed the system as strong in structure but in need of refinement in content granularity and contextual accuracy.
-E.6 Sepsis
-Experts agreed with most alignment scores, particularly those labeled partial, since the claims captured correct but incomplete aspects of the clinical reasoning process.
-For instance, “One risk factor for sepsis is advanced age” was appropriately labeled as partial under “Elderly Susceptibility,” as it was factually true but lacked the specific clinical threshold (age $\geq$ 65).
-Similarly, claims about single vital sign abnormalities were correctly labeled as partial since sepsis criteria like SIRS require multiple indicators.
-However, experts disagreed when the model evaluated isolated claims that should have been considered in combination.
-Statements like “A high triage temperature indicates fever” and “An elevated respiratory rate is another risk factor” were accurate but insufficient alone to establish “SIRS Positivity,” leading to misalignment between formal criteria and contextual reasoning.
-Additionally, experts highlighted that the LLM would generate explanations that involved risk factors that are not the same as diagnostic criteria for sepsis — while risk factors increase the likelihood of sepsis, there is no causal relationship.
-At a higher level, experts appreciated the clarity of claims tied to objective measures such as qSOFA scores and age but highlighted a persistent tension between rule-based alignment and real-world clinical reasoning.
-They emphasized that diagnostic logic in practice is more flexible than rigid scoring systems, exposing a conceptual gap that the model did not fully capture.